### v1

# Euclid Quick Data Release (Q1)

Euclid Quick Data Release (Q1)
Light emission from galaxies exhibit diverse brightness profiles, influenced by factors such as galaxy type, structural features and interactions with other galaxies. Elliptical galaxies feature more uniform light distributions, while spiral and irregular galaxies have complex, varied light profiles due to their structural heterogeneity and star-forming activity. In addition, galaxies with an active galactic nucleus (AGN) feature intense, concentrated emission from gas accretion around supermassive black holes, superimposed on regular galactic light, while quasi-stellar objects (QSO) are the extreme case of the AGN emission dominating the galaxy. The challenge of identifying AGN and QSO has been discussed many times in the literature, often requiring multi-wavelength observations. This paper introduces a novel approach to identify AGN and QSO from a single image. Diffusion models have been recently developed in the machine-learning literature to generate realistic-looking images of everyday objects. Utilising the spatial resolving power of the Euclid VIS images, we created a diffusion model trained on one million sources, without using any source pre-selection or labels. The model learns to reconstruct light distributions of normal galaxies, since the population is dominated by them. We condition the prediction of the central light distribution by masking the central few pixels of each source and reconstruct the light according to the diffusion model. We further use this prediction to identify sources that deviate from this profile by examining the reconstruction error of the few central pixels regenerated in each source’s core. Our approach, solely using VIS imaging, features high completeness compared to traditional methods of AGN and QSO selection, including optical, near-infrared, mid-infrared, and X-rays. Our study offers practical insights for refining diffusion models and broadening their applications throughout the Euclid survey area, underscoring the utility of this approach in diverse astronomical contexts beyond just AGN identification.
Key Words.:
Galaxies: active – techniques: image processing – methods: data analysis – methods: observational1 Introduction
Ever since the discovery that the mass or dynamics of a spheroid within a galaxy correlates with the mass of any detectable supermassive black hole (SMBH) at its centre (e.g., Dressler 1989; Kormendy & Richstone 1995; Magorrian et al. 1998; Ferrarese & Merritt 2000, and more recently, Sahu et al. 2019; Davis et al. 2018, 2019 and references therein), it has been believed that most – if not all – galaxies with a spheroidal (or pseudo-spheroidal) component contain a SMBH at their centre. If galaxy’s emission shows evidence for its central SMBH being actively fuelled, it will emit non-thermal radiation leading it to be classified as an active galactic nucleus (AGN). Depending upon the luminosity of that emission and its contrast with the stellar emission from the surrounding galaxy, it could have a classification ranging from a Seyfert galaxy (for the weakest emission) through to a quasi-stellar object or QSO (for the most luminous).
For imaging in wavebands such as the optical and near-IR, the presence and detectability of emission from the area directly around a SMBH depends on multiple factors. This can include the rate of infall of material into the region around the SMBH, the amount of obscuration (absorption or scattering) towards that region, the contrast between the SMBH-related emission and that of the surrounding galaxy, the redshift, and the intrinsic compactness of the stellar distribution of the host galaxy. The growth and mass of a central SMBH is intimately related to the evolution of its host galaxy, not least through the action of energy, momentum and radiative feedback from outflows originating from the area around the SMBH on the properties of the stellar populations and ISM (see Harrison & Ramos Almeida 2024 for a recent review). Consequently, identifying and characterising SMBHs, particularly during periods of active feedback is central to the study of SMBH and galaxy evolution.
A more precise classification of a source as an AGN is possible by using a much wider range of multi-wavelength data. This can include spectroscopy and considering the much broader spectral energy distribution of the galaxy and/or the candidate AGN. However, it is still worth considering the efficient selection of AGN candidates from single-band imaging as a first step in identifying large samples of AGN for further study. In particular, with the advent of the uniform, deep and high spatial-resolution optical imaging data set from the Euclid mission, covering a large fraction of the extragalactic sky, there is the potential to derive unprecedentedly large samples of AGN across a wide range of (optical) luminosity and redshift parameter space with the use of an appropriate and sufficiently efficient technique. The technique can then be subsequently refined by feeding back the result of using other multi-wavelength data, allowing us to confirm or reject candidates. With sufficient refinement, this further assessment step subsequently becomes either unnecessary or at the very least an efficient stage in compiling reliable samples. This then opens the door to production of unprecedentedly large and reliable samples of optically-selected AGN.
In this work, we present a technique that can achieve the above goal. In the following we describe the use of diffusion models for generating samples of AGN candidates using Euclid VIS band (Euclid Collaboration: Mellier et al. 2024; Euclid Collaboration: Cropper et al. 2024) imaging alone. Diffusion-based methods are machine-learning models used to generate or reconstruct data by evolving a random distribution of pixels into a structured output over several steps. This technique outperforms traditional methods like median pixel substitution or interpolation, which often struggle to recreate complex structures within images. By using these models for ‘inpainting’, where we adaptively recreate parts of the image, we can measure the discrepancies between the original and generated pixels. Examining these errors allows us to separate AGN, QSOs, and other sources from the broader galaxy population.
Briefly, the technique takes an image of a galaxy that may or may not have AGN emission at its centre and excludes those central pixels that may be influenced by AGN emission. The model then predicts (‘inpaints’) the underlying stellar light profile in those pixels from a comparison of the rest of the galaxy image to a model derived from a very large sample of similar galaxy images. Any significant difference between the prediction and true image in those central regions indicates the potential presence of AGN emission. Because AGN are comparatively rare, the sample of galaxy images used in the generation of the model does not need to be filtered to reject AGN, the method will be inherently biased to the ‘typical’ behaviour of galaxy light distributions and so naturally down-weights or excludes predicting central regions with AGN emission.
While this paper details the basic technique applied to VIS imaging data, as this is a machine-learning technique, future enhancements would allow incorporation of the results of a detailed assessment step, involving other data sets, in order to further improve the accuracy and reliability of the AGN candidate samples produced directly by the technique. Such steps could include multi-wavelength cross-matching or spectroscopic validation, from which we can ultimately reduce the need for further time-expensive refinement steps. Additionally, the technique as presented here uses a standard treatment of diffusion models which in principle can be significantly further optimised for the specific characteristics of the Euclid data set, thereby improving further the efficiency of the processing of the Euclid data. Such work would be the subject of future papers.
2 Diffusion-based inpainting
2.1 Generative models
Generative models such as generative adversarial networks (GANs, Goodfellow et al. 2014, 2020) or variational autoencoders (VAEs, Kingma 2013) are trained with the aim to learn the underlying distribution of a data set to generate novel examples.
Using a pair of competing networks, GANs ensure that once trained the generated images are of high quality, leading to the creation of many artificially generated data sets that are indistinguishable from the original data set (Karras et al. 2020). However, the adversarial training loop is often difficult to train, often leading to mode collapse, meaning the generator focuses on only a small subset of the total search space leading to generating samples of similar instances (Thanh-Tung & Tran 2020).
On the other hand, VAEs use an encoder network to compress the data into a latent space of fewer dimensions. The network then uses the latent space to recreate the input through an often symmetric decoder network. VAEs provide better coverage over the search space, improving stability during training as well as increased variety in generated outputs. However, the drawback for VAEs becomes apparent when handling more complex, high dimensional data. The use of latent spaces does not guarantee a lossless compression of information, often leading to accurate but blurry outputs.
2.2 Diffusion models
Diffusion-based models provide an alternative process for image generation that allows for both good coverage, as well as the ability to represent complex data distributions (Dhariwal & Nichol 2021). For both GANs and VAEs, their significant advantage with respect to diffusion models is that they offer faster sample generation and lower computational costs. Whereas diffusion models trade off speed for superior sample quality, mode coverage, and flexibility in controllable generation making them our preferred choice in the task of AGN identification.
This paper makes use of one such implementation, denoising diffusion probabilistic models (DDPMs, Ho et al. 2020). Diffusion works on the following premise: by continually adding a small amount of noise to an image, eventually one is left with an image of complete noise without any remnants of the original input. If the noise is generated in a stochastic but consistent way, such as from a Gaussian distribution, we can use a network to learn the dynamics of the noise and perform a reverse process. By predicting the noise that was added to an image at timestep , we can remove this noise to produce image . This forms an iterative process originating from some final timestep , an image of pure noise, to a realistic image from the data distribution at timestep .
The formal definition of this diffusion process, following Ho et al. (2020) and Nichol & Dhariwal (2021), is expressed as which represents the joint distribution of a sequence where each image is progressively noised.
where, specifies the conditional distribution at each timestep , modelling the incremental addition of noise:
The term defines the variance of the Gaussian noise added at each timestep. By ensuring a gradual increase in noise, the model is more easily able to learn the transition from the data distribution to the noise distribution.
With a large enough and an adjusted time-dependent variance , will approach a Gaussian distribution that is isotropic, meaning uniform in all directions. For the denoising process, starting at a , accurately modelling is not tractable, requiring a neural network to approximate it. An example of the full diffusion process can be seen in the top panel of Fig. 1.
Smith et al. (2022) also utilised the DDPM framework to generate a data set of realistic galaxies. The model is trained using the Photometry and Rotation curve OBservations from Extragalactic Surveys (PROBES) data set (Stone & Courteau 2019; Stone et al. 2021), a collection of large, well resolved objects that feature significant internal structure. They provide analysis of the quality of their generated images, showcasing the diffusion model’s ability to create not only visually realistic images, but also how they feature similar physical property distributions such as half-light radius and flux-space colour values. They also briefly explore the use of inpainting with their model to remove satellite trails.
When researching diffusion models, papers utilising score-based models are often used and spoken of interchangeably, as they are different implementations of the same generative process (Smith et al. 2022). Rather than the fixed sequence of timesteps to denoise data, score-based methods apply stochastic differential equations (SDEs) to estimate the data distribution’s gradient (score). Working with SDEs and gradients allows for a continuous range of possible diffusion paths. It is for this reason that score-based models can be classed as the more general framework for diffusion-based generative models. The use of score-based generative models within astronomical applications include galaxy image deconvolution (Adam et al. 2023; Spagnoletti et al. 2024), gravitation lensing analysis (Adam et al. 2022; Remy et al. 2023), and deblending (Sampson et al. 2024).
2.3 Noise scheduling
The larger the number of total timesteps , the more refined the model can become at accurately generating realistic images, since less complex noise dynamics are required to be learnt for each timestep . This does, however, require the model to learn the dynamics of more timesteps, leading to extra overhead in both training and inference, since more steps must be performed to traverse from pure noise to realistic images.
In the original DDPM paper, Ho et al. (2020) made use of a linear schedule for the amount of noise added to the image at each . Although the produced images were competing with state-of-the-art methods, Nichol & Dhariwal (2021) found that a linear schedule was not well suited to low resolution images and proposed using a cosine-beta noise schedule as a replacement. By stretching the noise levels so that less noise is added in early iterations and more in the later, the process overcame the issue of images became too noisy too quickly. Figure 2 shows how the scheduler impacts the parameter used in Eq. 2. Equation 3 shows how is calculated, where the reader is directed to Nichol & Dhariwal (2021) for specifics on its derivation:
where
All models in this paper make use of the cosine-beta scheduler. The impact of this scheduler on the high dynamic range nature of astronomy data is explored in detail in Sect. 5.
2.4 Conditioning the model
Using the pipeline of DDPMs allows us to impose conditioning on our input space. Throughout training, the model is already conditioned on . Due to the cosine scheduler discussed previously, the model will alter an image very differently at a timestep close to , where the image is still very noisy, than when the image is close to timestep and near completion. However, we are not limited to only conditioning by timestep, often multiple parameters are provided alongside each image. Typically, this is in respect to ’class-conditioning’, where users can tune the expected output of the diffusion model to a specific class or category (Radford et al. 2021; Zhang et al. 2023). By providing an embedding space during training, the model learns not only the dynamics of the noise but the inherent differences of noise and non-noise pixel values for each of the representations. This conditioning provides small nudges to the random walk, guiding the denoising to the respective population of the class. Providing rich representations to the model has led to the impressive text-to-image capabilities of popular diffusion-based implementations, such as Dall-E (Ramesh et al. 2022) or Imagen (Saharia et al. 2022).
This type of conditioning allows the user to direct the image generation to a particular class or style of image. However, with the possibility of complex prompts and classes, it may be difficult to know exactly what the output image will look like. Furthermore, with diffusion’s ability to cover large areas of the search space, a repeat sample of the same prompt is likely to produce an image very different from the first, even if the generated image belongs to the requested class. This is due to the one-to-many relationship between classes/prompts and their subsequent generated outputs.
In a situation where the user requires consistency in part of the image, a more direct manipulation of the produced pixels is required. A process known as ’inpainting’ (Pathak et al. 2016; Yeh et al. 2017; Yu et al. 2018; Wang et al. 2018), forces conditioning of specific pixels within the image, allowing the model to fix certain parts of the image whilst generating others – all whilst remaining coherent and consistent across the image as a whole. Allowing the user to selectively remove and replace certain aspects of images has gained a lot of popularity in the last few years due to its mainstream adoption by nearly all smartphone brands (Sargsyan et al. 2023).
2.5 Repainting
This paper uses a specific implementation of inpainting using diffusion known as Repaint (Lugmayr et al. 2022). Due to the model being conditioned on timestep , where it is expecting the whole image to be at a certain noise level, we can not simply take an image and a corresponding mask and replace the masked pixels with noise. The model itself does not know that a mask has been applied nor does it know that the inverse mask should remain unaffected. Instead, it will continue to remove noise across the whole image as usual, leading to the clean parts of the image becoming corrupted. It is also not possible to generate an entirely new image and copy over the masked pixels to the original image as there will be no continuity between the existing and new elements.
To combine the parts of the image we want unchanged, and the parts we want to generate, both the forward pass and the backward pass of the diffusion process must be utilised. This pipeline needs to generate an image that incorporates the mask and the fixed pixels for each timestep , as shown in Fig. 1.
Input
The main input of each timestep of the repaint pipeline is the denoised image, , from the previous timestep. To aid in the understanding of the pipeline, this image is referred to as for this discussion. At the first iteration when , the input is simply an image of pure Gaussian noise.
Fixed pixels
To ensure that the pixels we want unchanged are still present at the end of the diffusion process, we must interject these fixed pixels during each timestep. However, as the model is expecting all pixels to be at a specific noise level, as described in Eq. 2, we must utilise the same scheduler used in training to add the correct amount of noise for the current timestep . Due to the stochasticity present in adding Gaussian noise to the image, the pipeline allows us to create different versions of the input with the required noise. By averaging these samples, the important details of the image are more consistently preserved, ensuring that once all iterations are complete, the fixed pixels look identical to the original input. We refer to the image formed from the average of the N samples as .
Utilising the mask
To combine the parts of the image we want to keep and the parts we want to generate, we utilise a mask , which is an array of equal size to our image, with pixel values of either zero or one. Pixels valued as one will be generated from scratch by the diffusion process, with pixels valued zero being replaced at each noise step with the fixed pixels in from the process described above. The two parts can then be combined into a coherent image and then passed back into the diffusion model to be further denoised and progress onto timestep ,
By denoising the combination of pixels we want to keep with the generated pixels, we condition the diffusion process to ensure that in the final image, all pixels are coherent, being conditioned on the details found within the fixed pixels.
Incorporating the Repaint method for AGN identification has the following benefits
Diffusion methods have been proven their ability in many applications to accurately recreate complex data sets, allowing for realistic generation of galaxy morphology images.
There are no constraints on the size and shape of masks used and, since Repainting is only applied at inference, no additional retraining is required.
There exists a large imbalance of normal galaxies compared to AGN, allowing for the model to learn a bias. This can be exploited to create an outlier detection-based classifier without explicitly requiring labels during training.
3 Data
3.1 Sample selection
Covering 63.1 deg2, the Euclid Q1 Data Release (Euclid Quick Release Q1 2025) contains just under 30 million catalogued sources (Euclid Collaboration: Romelli et al. 2025) in the Euclid Deep Fields North (EDF-N), South (EDF-S), and Fornax (EDF-F). In this work we use exclusively the VIS images (Euclid band), because we aim to identify AGN candidates on a single image. However, due to the varying distances and sizes of sources, an appropriate subset must be chosen to ensure the model is presented with a sufficiently broad selection of resolved galaxies. The following criteria are applied to the 30 million source catalogue of Q1:
the source must be detected in the VIS band (vis_det_euclid0);
the source must have a magnitude brighter than (23.5 2.5log(flux_detection_total_euclid) 22.5);
the source must not have any defects or abnormalities as signified by the detection flags (det_quality_flag_euclid = 0 and flag_vis_euclid = 0);
the source must have low probability of being spurious (spurious_prob_euclid 0.2);
the source must have a large segmentation area (segmentation_area_euclid200);
the signal-to-noise ratio of the total flux measurement must be sufficiently high ( 15); and
if the signal-to-noise criterion is not met, then the segmentation area of the source must be very large (segmentation_area_euclid 1200), as defined in Euclid Collaboration: Walmsley et al. (2025).
After this sample selection we are left with 1 142 606 sources across the three Q1 fields. VIS image cutouts (6464 pixels) were created for each of the respective sources, where problematic images, such as those at the very edge of tiles where a full cutout could not be created, were removed. This provided us with images and catalogue data for 1 060 864 sources. We then apply a 80 : 20 train and test split. Even though the inpainting and its results are a separate pipeline to the diffusion training, to prevent any potential data leakage, all of the decision boundaries created are based solely on the training set. When showing classifier performance, all results are from the unseen 20% of the data. This is to prevent any potential memorisation of the images from training impacting the results. However, in our analysis, no obvious differences have been found in the inpainting results between the two sets, showing that the model is not simply memorising the training set.
To explore the impact point-like sources have on training and the inpainting process, a second smaller selection was created that combined all the selections above with the following:
the source must have a low probability of being star. (phz_star_prob_phz_class 0.1); and
the source must have a low probability of being point-like. (point_like_prob_euclid 0.2).
After removing problematic images as before, this smaller data set contained 794 624 sources.
3.2 Comparisons with other classifications
Although the training of the diffusion model does not make use of any labels, it is important to compare the distributions of metrics over different source selection criteria. This allows us to compare both the performance of detecting particular sources but also to determine what sources or factors may hinder the pipeline’s ability to classify objects sufficiently.
In particular, we use some of the sample definitions discussed in Euclid Collaboration: Matamoro Zatarain et al. (2025). A selection of stars are used from DESI (DESI Collaboration et al. 2024) and Gaia (Gaia Collaboration: Vallenari et al. 2023; Gaia Collaboration: Bailer-Jones et al. 2023). The normal galaxies sample is also from DESI (DESI Collaboration et al. 2024). Comparisons are made to selections of AGN and quasars using Euclid photometry with the _ and _ selections from Euclid Collaboration: Matamoro Zatarain et al. (2025) and the 2-colour selection from Euclid Collaboration: Bisigello et al. (2024, hereafter Bisigello24-A) . Additional AGN and QSO selections are used from DESI (DESI Collaboration et al. 2024), Gaia/Quaia (Storey-Fisher et al. 2024), as well as the 90% reliability WISE sample from (Assef et al. 2018), which is referred to as R90 for the remainder of the paper.
As is particularly true for AGN and QSO, it is impossible to create a selection criterion that will identify both with high completeness and high purity a single population of objects. Instead, we will use the aforementioned samples to assess contamination and purity of our method to the extent that is possible.
3.3 Data expectations
Although, by definition, a difference between a galaxy with and without a bright central component should be found, it is important to validate that this behaviour can be measured with the Euclid VIS images. To quantify this difference we analyse the relationship between the brightest central pixel and its surrounding pixels. By producing ratios comparing the central pixel to 1 and 2 pixels wide regions around the central pixel, we form a rate of change (ROC) metric over sources, as described by
where
and and are the mean values of the surrounding rings of 1 (total 8 pixels) and 2 pixels (total 16 pixels), respectively. is the pixel value of the brightest pixel within the centre 99 pixels.
Pixels surrounding the brightest pixel, which share similar values to the maximum will produce a ROC value closer to zero, implying that the source does not feature a prominent active component. Sources featuring a single very bright pixel, surrounded by fainter pixels will have a ROC value closer to its maximum of 2. The produced distributions of sources, analysed according to both magnitude and various selection criteria (Sect. 3.2), are shown in Fig. 3. It is evident that sources with a high flux point-spread function (PSF) light profile show the highest ROC values, as expected. This effect is seen in the star sample and the QSO dominated selections (R90 WISE, Quaia, DESI-QSO, and Euclid-QSO). On the other hand, spectroscopically confirmed normal galaxies show very broad ROC distributions.
3.4 Addressing the use of asinh for image scaling
| Galaxy Overlap | Without asinh | With asinh |
| AGN | 0.75% | 1.15% |
| QSO | 0.20% | 7.82% |
| Stars | 0.17% | 1.86% |
Given its prevalence in galaxy morphology-based research (Lupton et al. 2004), the effect of using an transformation on the data is also investigated. The function, which is often to the benefit of visual inspections, reduces the differences between the brightest and faintest pixels. However, when attempting to differentiate galaxy and non-galaxy sources, the function becomes a hindrance, causing the distributions to more significantly overlap, as shown in Fig. 4. To generate a measure for class overlap we measure the ROC for each image, taking the mean of each non-galaxy class. We use the percentage of our true galaxy sample that have a ROC value over the respective class mean as our overlap metric. The extent of the overlap is showcased in Table 1, where galaxy overlap increases by at least a factor of 2.5 and in the extreme case, nearly 50 times more galaxies overlap with stars. In the following sections of the paper we will discuss the set up, training and performance of the diffusion model that automatically learn to reconstruct the central emission expected for the normal galaxy population, and thus by contrast allow us to identify stars and AGN.
4 Training and inference pipeline
The diffusion and inpainting methods are two distinct pipelines, one for the initial training of the model and a second for the actual inpainting. Training the diffusion model makes use of the whole image and learns to predict what noise to remove at each timestep. This allows for the generation of novel examples from the trained distribution.
4.1 Training the diffusion model
For training, we use 6464 pixel VIS cutouts for each of the selected sources discussed in Sect. 3.1. We use the training pipeline described in Dhariwal & Nichol (2021)111Available at https://github.com/openai/guided-diffusion with the loss function
The hybrid loss is a combination of the mean squared error (MSE) and the variational lower bound (VLB). Using MSE ensures the output pixels of the model are numerically similar to the original image, maintaining structural and visual consistency within the generated images.
The first term in 6 is
where and are the true and predicted noise added for a timestep, respectively.
The VLB loss, shown in Eqs. 8 to 10, measures how well the diffusion process has modelled the underlying distribution of the image data, ensuring that the image as a whole is coherent and plausible for the data. The VLB loss has a weighting applied so as to not overpower the MSE loss:
By utilising the Kullback-Leibler (KL) divergence (Kullback & Leibler 1951), a measure of the difference of two distributions, the loss acts as a regulariser to prevent simple memorisation and instead optimising for learning general features and patterns present in real data.
By combining both losses, the model becomes accurate at generating realistic data at both the micro (pixel-wise) and macro (image-wise) levels. This is vital in ensuring the model learns the dynamics of morphology images through the diffusion process, allowing it to effectively recreate essential features through repainting. Throughout the rest of the paper the loss presented in Eq. 6 is referred to as the default loss.
Alternative loss
Making use of standard metrics whose primary use-case is data on a consistent scale, such as 0–255 pixels, may not function optimally with the high dynamic range of VIS data. Specifically, since MSE is not scale invariant, having images that feature pixels that are an order of magnitude brighter, results in relative errors increasing by a square of that difference. This is typically a benefit as it forces outlier predictions to be severely penalised. However, with our data, this results in brighter images being disproportionately penalised compared to fainter ones, even if the predictions are both the same relative distance away from the true prediction.
To address this, we applied a normalisation to by dividing each image’s error by its respective maximum pixel value, which could be from a brighter external source. This allows for the errors of both bright and faint images to scaled more equitably. The output of this loss is still combined with errors calculated in . Throughout the remainder of the paper, we refer to this variant of the loss as the normalised loss.
To visualise the impact of amending the loss function, with respect to an images brightness, Fig. 5 shows how the error from the unmodified MSE changes as the model is trained for longer. Each line represents the error for a collection of images in a random batch, with the increased transparency indicating the passage of time where the model had been trained on more epochs. The final epoch of each model is presented by a black line. The effect of removing point-like sources, as discussed in Sect. 3.1, is also explored.
An initial observation that can be made with this representation is the distribution of the number of images of a particular brightness. By focusing solely on the black lines, where the impact is most noticeable, each vertex on the line represents an individual image. The more frequent updates the line has within a range of maximum pixel values, the more images in our data set exist within that region. As the line has more updates when the max pixel value is less than one, there are significantly more images with fainter pixels. This further highlights the disproportionate penalisation of MSE as it assigns the majority of the error to a minority of the data.
Although all four models arrive at a similar minimum loss, and begin the training process near-identically, their respective behaviour throughout training differs significantly. For the full data set with the default loss (top left of Fig. 5), training prioritises brighter images, leading to some significantly low errors towards the rightmost data. However, this comes at the expense of fainter data, where wild fluctuations in error show that the model’s output is not very robust when it comes to precise predictions. This happens even when MSE might suggest low error, since pixel values below 1.0 result in minimal losses despite the predictions being relatively incorrect.
For the normalised loss with the full data set (top right), since the penalisation is shared more fairly across pixel values, it is the density of images that has the most influence. Early on in training we can see that as the majority of data reduces in error, the brightest images remain at near-starting performance. This continues until the errors of the fainter images become so low that model is forced to prioritise the bright images. The key difference here is this only happens after the rest of the images are at a sufficient standard. The final outcome is fainter images performing well and the entire scale of images having the same maximum error.
The two models which feature the reduced training set (bottom row) share similar dynamics during training, regardless of the loss used. This is largely due to many point-like sources also being some of the brightest images and therefore the scale of the data set is reduced. As the most problematic data are removed, the impact of disproportionate penalisation is significantly reduced, allowing the fainter sources to be the focus during training. After analysing the results from each of the four models, which will be discussed in Sect. 6.2, we chose to use the model trained on the full data set using the normalised loss.
4.2 Mask creation
Due to the nature of the application, our aim is to mask out any suspected AGN without the need to rely on labels ahead of time. Each mask is positioned so that brightest pixel in the centre of the source becomes the centre of the mask. Although each image is centred on the source, various interactions and dynamics allow for suspected AGN to be offset from the centre. To allow for some deviation, a 99 pixel window around the centre of the image is used to detect the brightest pixel. This offset allows for the correct masking of less symmetric sources whilst limiting the likelihood of masking an adjacent source by mistake. The masks for each image are assigned dynamically throughout the data loading pipeline.
Throughout this paper we use a 55 pixel mask, unless explicitly stated otherwise, centred on the calculated brightest pixel. A 55 pixel mask, which covers 0505, allows us to effectively cover the AGN contribution whilst minimising loss of the surrounding galaxy structure. Given that the VIS PSF has a full width half maximum (FWHM) of 013 (Euclid Collaboration: Mellier et al. 2024), the AGN’s light is spread over multiple pixels. Taking into account the redshift of the galaxy, AGN within nearby sources () will cover 4.8 pixels, ensuring that we can fully encapsulate the core without unnecessary overlap of the host morphology. By retaining as much as the host galaxy information as possible for inpainting, we also reduce the risk of introducing artificial features into the reconstructed image.
4.3 Reconstruction rescaling
An important difference with the processing of data in this work compared to other morphology-based papers is how the data input into the network are the raw values direct from the sensor, rather than images binned to the standard 0–255 pixel scale range. As comparisons of individual pixels are being made on such a small mask, the relative differences become much more impactful and would likely be lost if values were binned.
One potential issue with this setup is that the libraries used for training are optimised for images within the standard pixel range. Although this did not prevent the model from effective learning the distribution, it does require a rescaling of the generated output as the range is restricted to between 1 and 1 to improve training stability.
We experimented with simply normalising the input data so that all our images were in the range 0–1; however, due to the large dynamic range of the data, the faintest pixels ended up being . Given the previous discussions on how MSE does not penalise smaller errors effectively, the model could not accurately recreate sources, outputting a near empty image of very low pixel values.
Instead, to allow the comparison of the original and outputted pixel values, we make use of the fact that 99.5% of the image (4071 of 4096 pixels when using a 55 pixel mask) are shared between the input and output. The following scaling can be dynamically applied using each images respective values to allow the non-masked pixels to be consistent:
where
and
By ensuring that the shared pixels are consistent, the inpainted pixels can also be correctly adjusted.
5 Background noise versus diffusion noise
It is important for clarification to address the distinctions between the definition and use of ‘noise’ in diffusion models compared to the traditional understanding of noise in astronomy-based data. In the context of astronomical imaging, noise refers to the random fluctuations inherent in observational data, often caused by e.g. background emission, electronics read-out noise, and atmospheric disturbances. This ‘background noise’ reduces the signal-to-noise ratio (S/N) of the image, affecting the ability to detect subtle details in any analysis.
In contrast, the diffusion process uses an intentional integration of Gaussian noise to facilitate training of the generative model. The pixels of image , follow the distribution: , rather than following the distribution of the background noise. It is unclear whether using the background noise as the pure noise sample would cause issues for the model. Given the aim of the diffusion process is to denoise a noisy image so that a realistic one emerges, the original noise should be largely, if not entirely, removed. Given the background noise constitutes any area that is not significantly affected by a source, the model may attempt to remove it which would create an unrealistic image. On the other hand, it may result in less pixels needing to be denoised, optimising the models ability at learning and processing the source features. Such an experiment is left for future exploration.
5.1 The effect of signal-to-noise
In machine-learning research, the improvements of generative models are often demonstrated with data sets that are of high resolution and typically noise free. This provides a level of consistency across the full data set, ensuring the benchmarks reflect the model’s performance under ideal conditions. Whether through loss functions (as discussed in Sect. 4) or in the practical application of the pipelines, how the model behaves on one part of the data will be similar to another.
However, the difficulty in learning the dynamics of our data are not consistent throughout the images. Higher signal-to-noise ratios and lower magnitudes provide higher contrast between the source features and the background noise, enabling the model to discern these features more accurately and faster than the low S/N counterparts. In the context of diffusion models, where the aim throughout training is to learn the dynamics of noise at different timesteps, the varying S/N of the images could again cause inconsistencies during training.
Figure 6 explores how this variability in data impacts the scheduled noise added to the data. By measuring the specific pixels that are dominantly background noise (top) and those that are significantly affect by the source (bottom), and further splitting images into S/N bins, we show the transformation from raw pixels to Gaussian noise across the entire scale of our data.
Analysing the top row, we can see that low S/N images are more significantly affected by the introduction of noise. This also has an effect much earlier on in the scheduler. The background noise converges to the diffusion’s pure Gaussian distribution, across all S/N bins, at between 1/2 () and 3/4 () of the total noise steps. When focusing on the pixels dominated by the source (bottom row), high S/N images are fairly unaffected for large parts of the schedule, with the brightest images not changing significantly until half way (). It is not until the scheduler is within the final 250 steps that the brightest images converge to the Gaussian distribution.
To visualise the true effect of this differing behaviour across S/N, Fig. 7 shows a sample of images throughout different timesteps. Each image shown is a sample from the respective S/N bins below. It is clear that the noise scheduling does not perform uniformly across the images.
To increase performance of a diffusion model, whether to recreate a data set more accurately or increase the complexity of the data you are modelling, the simplest way is to increase the number of unique denoising steps performed. The important part here is the requirement to be unique, as simply adding more timesteps often results in the additional timesteps (especially in the latter stages) looking too similar and therefore become just additional overhead. This is the exact reason Nichol & Dhariwal (2021) implemented the cosine-beta scheduler, discussed Sect. 2.3, the original linear scheduler produced images that were pure noise too early into the process.
The images in Fig. 7 show that there are further optimisations that could be made to improve the scheduler for our data set. For optimal performance, one should expect the entire grid of images to behave like those in final 2 or 3 columns. A scheduler that adapts to the specific S/N of the image would be required and is left for future work. As our work is specifically using the reconstruction loss of inpainting as a classifier, the current scheduler is sufficient for our science goals.
6 Results
To explore the ability of the model at recreating galaxy images, we apply a variety of different masks to a selection of unseen images. Figure 8 shows the original image on the leftmost column, followed by the regenerated images in subsequent columns. With each mask, shown on the top row, the white area shows the pixels that are fixed from the original image whereas the black areas are generated using the diffusion model.
As our model’s aim is not to generate large collections of synthetic data, the variety and scientific validity of the produced galaxies are not put under high scrutiny, instead we only check that there are no obvious catastrophic mistakes. In the last two columns, we can see that due to the large mask and fewer pixels to guide the reconstruction, some of the images produce a visible boundary around the image where the combination of fixed pixels and generated pixels are not as seamless as the other images. Given the masks used for AGN detection are significantly smaller than those presented in Fig. 8, the fraction of pixels that the inpainting is conditioned on will be much higher, providing tighter constraints on the output compared to these examples.
6.1 Inpainting of AGN
For our inpainting, we use a 55 pixel mask centred on the brightest pixel, following the constraints described in Sect. 4.2. Comparing the expected pixel values with the generated outputs, Fig. 9 shows the model’s performance. The assumption that the model will predict a fainter core than the original does appear to be true, with a near global bias for underpredicting as shown by the dashed line. The majority of sources behave in a consistent way with the exception of a smaller cluster of sources that predict more than an order of magnitude fainter than the original. This distinct cluster could provide sufficient candidates for AGN.
The horizontal bar in the leftmost panel of Fig. 9, where the maximum output value appears to equal one, is a side effect of Eq. 11. The values themselves are not exactly one, with 1% of the data fall within of having a maximum output of one. There does exist a convergence point in Eq. 11 , where when the maximum and minimum pixel values of an input image equal one, the output of the rescaling always equals one; however, such an image would have all its pixels valued one, and so is not present in the data. Although there is not a specific convergence point that explains this trend, the majority of the images that fall on this line occur when the network outputs a maximum pixel value close to one. However, there are many more images that have a network output close to one that are not close to this line, as well as images close to the line without a network output close to one. We have verified that both point-like and extended source appear within this line, therefore due to the low number of sources impacted, we leave more in-depth exploration to future work.
To further validate the output of the model and show it is producing realistic outputs, Fig. 19 shows how the ROC metric compares between input and output images. The significant overlap, especially with respect to peak of each distribution, shows that the model is accurately recreating the dynamics of the data.
To analyse how the error of each subset of class candidates have performed, comparisons are made between the original pixels and inpainted reconstruction for each image. The MSE of the pixel values are shown in Fig. 10. Each subset is binned according to its magnitude, providing a distinction between the relative brightness of each image.
There is a clear distribution difference between the respective class candidates. Stars perform the most consistently across magnitudes, producing the highest MSEs. There is an inherent trend for how high the error of the predictions are with the relative brightness of the classes, with stars and the brightest QSOs featuring the highest errors across all magnitudes. This is then followed by QSOs at higher magnitudes, and the lowest magnitude AGN candidates. Finally, the galaxy sample, along with the faintest AGN candidates produce much lower errors. Since these classes are candidates, it could also highlight the presence of contamination within the AGN labels. However, the large number of DESI-detected galaxies showing vastly reduced errors implies that the model does develop a bias for recreating galaxies with little or no AGN component.
The second important finding is the trend between VIS magnitude and error rate across all classes. The brighter a source is, the higher its recorded MSE. This matches the initial assumptions of the method, where, as the sources increase in magnitude, differences in brightness values across galaxies become harder to differentiate. This leads to fainter potential AGN candidates looking more similar to a normal galaxy, causing the inpainting to reconstruct more accurately and therefore less likely to be included in our selection. However, the relationship between brightness and error is once again more influenced by the scale-variant nature of MSE. However, the lack of significant overlap for many of the selections with the galaxy class still make this metric useful.
To showcase how the performance of the inpainting method impacts the overall image, Fig. 11 presents the input, output and residuals for a selection of sources, including a close up of the masked area both before and after inpainting. It is clear to see how bright central cores can be entirely removed with inpainting, being replaced by regions that are consistent with the surrounding pixels. The model is also able to recreate images that do not feature any significant core accurately. This shows the model has a good internal representation of galaxies and how they should behave.
6.2 Max pixel difference ratio
Although reconstruction error allows stars and a subset of AGN to stand out from the rest, a substantial overlap between suspected AGN and DESI-classified galaxies remains, especially at higher magnitudes. Even though measured on a pixel-by-pixel scale, MSE averages away specific pixel differences, leading to less nuanced differences between input and output. As an example, if the reconstructed image ended up being the correct values but just shifted by one pixel, this would be penalised heavily under MSE. From the perspective of this paper, such a scenario would imply that the model knew a bright pixel was valid and therefore, should not be penalised. Another example would be if only the brightest pixel is different in the reconstruction. MSE would significantly reduce the impact of this measurement, potentially scoring similar to a source that gave small errors in every pixel within the mask.
To ensure such cases are not overlooked, we measure the pixel value ratio between the brightest pixel from the input mask and the brightest pixel from the reconstruction. As shown in Fig. 12, the produced distributions show a much tighter spread of values for many of the AGN selections, with a skew towards higher scores. The majority of the galaxy subset produce a ratio less than 10, whereas a large population of suspected AGN from each selection feature a larger ratio.
Another important difference between this metric and MSE is how the results are relative, reducing the impact of magnitude on the recorded values. Many of the high density peaks remain consistent, even as the magnitude of the sources increase. This implies that this metric is more robust at detecting fainter AGN than MSE.
The distribution also highlights the scale of the extreme values with many sources showing pixel values approaching 100 times greater than their counterparts. Exploring these large disparities can provide insights into the types of sources that are significantly sensitive to the model’s capabilities. Sources producing the highest values for the ratio are those where the reconstruction produces a cutout that is significantly fainter than the original image. As can be seen in Fig. 20, this space is dominated by point-like objects. Having the sharp falloff of pixel intensity in such a small area leads to the mask covering the majority, if not the entirety of the source. When the model attempts to reconstruct the missing pixels, it generates values that are coherent with the surrounding pixels. In these cases, the best match for empty space being more empty space.
The behaviour of the previous examples match with the expected behaviour of the model and the initial assumptions of the paper. Applying inpainting to this task is expected to produce outputs that are similar to, or produce a flattened light profile of the original pixels. However, there exists a subset of sources that exhibit the opposite behaviour and generate a much brighter collection of pixels. Figure 21 shows examples of such sources. When analysing the sources that exhibit this behaviour we find that they are overwhelmingly low S/N images. Given the difference in behaviour of these images during training, as shown in Fig. 7, it is clear this has had an impact on the model’s behaviour with these types of images.
Because we would like to limit this behaviour, we analyse how many of the unseen sources produce this type of output. The number of sources whose max pixel difference ratio is less than one for each of the four models discussed in Sect. 4, are shown in Table 2. The normalised loss helps significantly in reducing the number of sources exhibiting this behaviour. The normalised loss trained on the full data set was selected as our applied model due to it further reducing the number of outlier predictions.
| Data set | Loss function | |
| Default | Normalised | |
| All data | 55 008 (6.49%) | 33 075 (3.90%) |
| No point-like | 52 313 (6.17%) | 39 507 (4.66%) |
6.3 Morphology-based performance
Given the variety of shapes and sizes of objects that will be witnessed using Euclid, it is important to verify that the model is able to handle the complexities of different galaxy morphologies. Utilising the refined morphology classifications provided from Euclid Collaboration: Walmsley et al. (2025), we provide examples of edge-on (disk-edge-on_yes_fraction ), spiral (has-spiral-arms_yes_fraction ), and mergers (merging_merger_fraction ) in Figs. 22, 23, LABEL:, and 24, respectively.
These images show how our diffusion model and the inpainting pipeline effectively adapt to different morphologies. For edge-on in particular, the region of bright pixels within the masked area is often narrower than the mask itself. Despite this, the inpainting correctly preserves the orientation of the features rather than uniformly filling the entire masked area with bright pixels.
6.4 Creating a classifier
Due to the difference in distributions between the two metrics, any non-overlapping cases can be found when used in combination. As seen in Fig. 13, there is a cluster of sources in the top right. The key aspect of this cluster is its clear boundary from the DESI galaxy selection. This is vital for ensuring high reliability in classifications. We create a simple decision boundary between the two clusters of points, as shown in Fig. 13. Although a more fine-tuned boundary would certainly be possible, especially when incorporating a further machine-learning model, it is important to show the benefits of utilising the inpainting method through only its error metrics. The equation of our linear boundary is .
6.5 Variance of reconstructions
Given the stochastic nature of the diffusion model pipeline, it is important to verify that any produced results can remain consistent if applied to the wider data set. Although the nonmasked region of the images will remain fixed over repeat samples, the random walk performed in the inpainting procedure could lead the samples to a different area of the search space. This would result in a different, but possibly equally plausible, output image.
We perform the inpainting pipeline seven separate times on the same subset of the data. To ensure that any low variance in samples is not caused by the model simply ‘remembering’ the true image from its training examples, all resampling is applied on images the model has never seen before. The mean and variance over a normalised by brightness MSE are shown in Fig. 14.
The top figure shows that irrespective of magnitude, the model achieves similar relative median error across all images, highlighting the impact of the normalised loss throughout training. However, the difference in the scheduler over images of different quality and brightness is made evident in the bottom figure. The consistency of generated outputs vary drastically as fainter (and therefore lower S/N) images produce widely different pixels when resampled.
6.6 Comparing to other criteria
It is important to compare our AGN and QSO predictions to more traditional methods that are also able to utilise only the image data. Given the assumption of our model that sources containing an AGN will feature a bright central component with a sharp falloff in the galaxy’s light profile, comparing to Srsic flux and point spread function (PSF) flux measurements is crucial because they capture the key concepts of our assumptions. The Srsic flux helps quantify galaxy light distribution and differentiate an AGN’s central brightness from the galaxy’s structural brightness. PSF flux isolates the galaxy’s core, capturing point-like behaviour. The distribution of sources with these parameters is shown in Fig. 15. The equation of this linear boundary is .
Our second comparison is with the original rate of change metric used in Sect. 3.3. As we were able to see a clear distinction between the galaxy class and the others, it is important to verify how such a metric performs as a classifier. This is especially true given such a metric can be generated from the raw image, with no machine-learning model required. The boundary for this classifier is .
The test set scores for each respective selection is shown in Fig. 16, with the collated results for stars, galaxies, AGN and QSOs shown in Fig. 25. All predictions are on images that were not used during training, nor used in the creation of any of the decision boundaries for the classifiers.
Given the variety and differences in the selection methods, there is an inherent incompleteness and potential for contamination, therefore our analysis focuses primarily on achieving sufficiently high recall. High recall ensures that our method is capable of identifying the broad range of AGN and QSO sources that are detectable through existing techniques. A high recall metric indicates that our classifier effectively captures the variety of sources each respective selection criterion is designed to detect.
While precision is also a consideration, our primary metric of success in this context is the diffusion model’s recall. Having the ability to recover a large fraction of known AGN and QSO types, across various selection methods, validates the robustness and comprehensiveness of our classifier. Due to differing constraints for each of the selections, we split the analysis to allow for a more representative comparison.
Galaxies
Our galaxy selection is determined by the DESI spectral type classification SPECTYPE=GALAXY (DESI Collaboration et al. 2024), with AGN contaminants removed following the procedure detailed in Euclid Collaboration: Matamoro Zatarain et al. (2025), hereafter called MZ25. Negative instances are formed from the remaining spectroscopically determined labels SPECTYPE=STAR and SPECTYPE=QSO (DESI Collaboration et al. 2024) along with the contaminants removed from the original SPECTYPE=GALAXY selections.
The predictions from each classifier for whether a source is a normal galaxy is simply the inverse of the boundary’s original predictions. As Fig. 16 shows, all classifiers capture nearly all galaxies across every magnitude bin. The precision scores also match extremely closely between classifiers, indicating similar contaminants proving difficult, especially for the brighter sources. Such contaminants can be seen in Figs. 13 and 15, where stars and DESI-selected QSO fall outside the boundaries.
The ability to separate normal galaxies from other extragalactic or stellar objects has benefits for both groups of data. Being able to create a pure galaxy sample is vital for the precise measurements needed for cosmology, as well as to further our understand of dark matter and dark energy, a primary object of the Euclid mission (Euclid Collaboration: Mellier et al. 2024).
Stars
The star selection is made up of the positive instances of Euclid selected (star_candidate_euclid, MZ25), DESI-selected (SPECTYPE=STAR, DESI Collaboration et al. 2024), and Gaia-selected stars.
Due to the assumptions made for this method, the incidental capture of stars within our classifications is an expected outcome. The near-perfect recall shows that our classifier is able to capture stars well. This performance is beneficial when attempting to capture a high purity galaxy sample as discussed in the previous section. However, as other works such as Euclid Collaboration: Matamoro Zatarain et al. (2025) have shown, there are metrics to mitigate the contamination of stars. This is shown in practise by the high precision, highlighting how we are able to isolate this star selection from other point-like sources.
However, a significant number of stars are likely to remain in our sample. Therefore, for all subsequent selections, we apply the cut phz_star_prob_phz_class, to predictions of all classifiers to minimise the contamination of stellar sources.
Euclid-detected QSOs
The Euclid photometric-based selections for QSOs include _ and _, as described in MZ25, and the Bisigello24-A 2-colour selection (Euclid Collaboration: Bisigello et al. 2024). This subset consider only sources with MUMAX_MINUS_MAG , a parameter that indicates point-likeness (Euclid Collaboration: Romelli et al. 2025).
For the _ and _ selections, we capture around 90% of the candidate QSOs. This coverage remains consistent across the available magnitudes. The performance compared to the other classifiers is similar, with ROC capturing a small percentage more sources, but not different enough to be significant.
The 2-colour selection proved harder to match for all classifiers, with recall dropping to 0.6 for the faintest objects. Again, no significant differences are seen between the classifiers.
DESI-selected QSOs
As these sources are assumed point-like, applying the above MUMAX_MINUS_MAG limit would be reasonable. However, due to the improvements in spatial resolution from DESI, at about , to Euclid’s now 02, some sources that appeared point-like through DESI, now appear extended. Therefore, to ensure all potential sources are found, we do not apply the restriction to this subset.
This selection provides the first evidence of significant differences within the candidates for each classifier. The inpainting classifier is able to capture a significant amount of the selection compared to the other classifiers. Throughout all magnitude bands, the PSF and Srsic flux classifier has worse recall. Whereas the diffusion model has higher coverage over the brightest sources (lowest magnitudes), the ROC classifer matched more candidates in the fainter sources.
Gaia-detected QSOs
For the Gaia-detected QSOs, we use the Quaia selection (Storey-Fisher et al. 2024). The subset consists of positive instances (qso_quaia, MZ25), with negative instances from Gaia-detected galaxies (in_galaxy_candidates_gaia, MZ25) and stars (star_candidate_gaia, MZ25). We once again apply the MUMAX_MINUS_MAG restriction to the sources.
After the initial magnitude bin, our classifier was able to match with a high proportion of the Quaia-selected QSOs. The other two classifiers achieved equal recall scores, but in the faintest detected sources our classifier was able to match with an additional 20% of sources.
The particularly poor precision for this selection across all the classifiers is due to the considerable imbalance between positive and negative predictions. With only 0.5% of the Gaia-detected sources being Quaia candidates, selecting more than just the candidates is likely, significantly impacting the precision.
AllWISE R90 AGN
Utilising WISE-detected sources allows us to verify our ability to select candidates that may have been obscured by dust. Using the selection criteria derived by Assef et al. (2018), we compare our candidates to the 90% reliability (R90) selection which is a high purity AGN sample.
The Assef et al. (2018) selection provides a more significant test for the classifiers due to no restrictions on point-like sources, as well as fewer contaminants that may have unintentionally improved the overlap of candidates. Similarly to the DESI-selected QSOs, our classifier shows a higher coverage than the other decision boundaries on the brightest sources. In the fainter sources, the diffusion and ROC classifiers are matched. Given the purity of the R90 selection, a large sample of suspected AGN remained unmatched by any of the classifiers, with the most matched magnitude bin achieving 80%.
Much like the Gaia-detected QSOs, the small fraction of candidates compared to the full sample of possible sources leads to many of the classifiers selecting sources outside the R90 selection. With a completeness of only 17% (Assef et al. 2018) there will be many mid-infrared AGN in the negative sample, which the classifiers are likely finding.
| Recall | Precision | AUC | |||||||||||
| Selection | 18 22.5 | Ours | PSF–S | ROC | Ours | PSF–S | ROC | Ours | PSF–S | ROC | |||
| Stars | 0.99 | 0.99 | 1.00 | 0.96 | 0.97 | 0.96 | 0.50 | 0.51 | 0.50 | ||||
| 0.97 | 0.98 | 1.00 | 0.84 | 0.85 | 0.84 | 0.51 | 0.53 | 0.50 | |||||
| Galaxy DESI | 1.00 | 1.00 | 1.00 | 0.74 | 0.73 | 0.73 | 0.64 | 0.63 | 0.64 | ||||
| 0.98 | 1.00 | 0.99 | 0.90 | 0.90 | 0.91 | 0.73 | 0.72 | 0.75 | |||||
| QSO DESI | 0.77 | 0.57 | 0.63 | 0.18 | 0.14 | 0.15 | 0.84 | 0.75 | 0.77 | ||||
| 0.90 | 0.84 | 0.96 | 0.59 | 0.67 | 0.62 | 0.93 | 0.91 | 0.96 | |||||
| QSO QUAIA | 1.00 | 0.98 | 0.98 | 0.01 | 0.01 | 0.01 | 0.59 | 0.58 | 0.58 | ||||
| 1.00 | 0.80 | 0.80 | 0.01 | 0.01 | 0.01 | 0.51 | 0.41 | 0.40 | |||||
| AGN Xray | 0.64 | 0.48 | 0.61 | 0.74 | 0.73 | 0.74 | 0.63 | 0.59 | 0.62 | ||||
| 0.51 | 0.38 | 0.53 | 0.86 | 0.86 | 0.87 | 0.53 | 0.51 | 0.55 | |||||
| AGN R90 AllWISE | 0.68 | 0.57 | 0.63 | 0.02 | 0.02 | 0.02 | 0.68 | 0.63 | 0.66 | ||||
| 0.59 | 0.45 | 0.60 | 0.03 | 0.02 | 0.03 | 0.75 | 0.68 | 0.75 | |||||
| QSO _ | 1.00 | 1.00 | 1.00 | 0.02 | 0.02 | 0.02 | 0.50 | 0.50 | 0.50 | ||||
| 0.94 | 0.92 | 1.00 | 0.05 | 0.05 | 0.05 | 0.49 | 0.47 | 0.50 | |||||
| QSO _ | 1.00 | 1.00 | 1.00 | 0.01 | 0.01 | 0.01 | 0.50 | 0.50 | 0.50 | ||||
| 0.95 | 0.91 | 1.00 | 0.04 | 0.04 | 0.04 | 0.50 | 0.46 | 0.50 | |||||
| QSO Bisigello24-A | 1.00 | 0.99 | 1.00 | 0.02 | 0.02 | 0.02 | 0.50 | 0.50 | 0.50 | ||||
| 0.96 | 0.97 | 1.00 | 0.03 | 0.03 | 0.03 | 0.50 | 0.50 | 0.50 | |||||
Xray-selected AGN
We compare our candidates to the X-ray selection from the work by Euclid Collaboration: Roster et al. (2025). The positive instances are when the calculated probability that the source is a galaxy is sufficiently low (Gal_proba_roster , MZ25). The negative instances are sources with Gal_proba_roster . Only sources with detection in X-rays are included in the results.
The main difference with the results of the X-ray-detected AGN is how the precision outperforms the recall across nearly every magnitude bin. This is due to low number of X-ray sources compared to the other selections. The recall performance between the diffusion model and the ROC classifier is very similar, with the PSF and Srsic flux classifier again struggling to match as many candidates as the other two.
Each of the selection criteria we compare with captures distinct subsets of the AGN and QSO populations, influenced by specific observational biases and methodological constraints. Therefore, it is only when collating these selections and by identifying trends and overlaps in the labels, that we can be confident that we sufficiently cover the spectrum of possible AGN.
Table 3 summarises the scores shown in Fig. 16. We combine the results into two magnitude bins 1820 and 2022.5, separating into brighter and fainter sources respectively. Due to similar scores in many of the selections, we highlight in bold any scores that outperform the other classifiers by atleast 0.05.
The results show that the diffusion method achieves high recall scores across all tested selections, most notably outperforming the other classifiers in the bright QSO DESI, faint QSO QUAIA and bright R90 selections. In other selections, it performs comparably to the ROC classifier, with both classifiers often significantly outperforming the versus Sérsi classifier.
6.7 Decomposition of the AGN Component
Given our model’s ability to classify suspected AGN, it is not unreasonable for the reader to assume that the residuals of the repainting process and the output image can independently represent the AGN and galaxy components. Although utilising inpainting for decomposition is a worthwhile future research direction, the validity of the residuals to accurately represent the AGN component for further scientific analysis have not been tested and as such cannot be used. More significant exploration of which mask to use will also be vital for such a use-case. A mask that adapts its size and shape for each input image will likely be necessary. For a deep-learning approach applying AGN decomposition on Euclid data, readers are referred to Euclid Collaboration: Margalef-Bentabol et al. (2025), hereafter called MB25, whose method provides AGN fraction () measurements that we use to analyse our AGN candidates.
The subset of Q1 used within this work and the subset tested in MB25, feature an overlap of 425 781 sources. The confusion matrix for AGN predictions between the diffusion classifier and sources with an (AGN candidate for MB25) are shown in Table 4. The majority of sources () are in agreement by both methods not to be AGN candidates. With the remaining sources, only are in agreement, with over five times more AGN candidates selected by MB25. When comparing all the AGN candidates selected using their respective subsets of data, this number reduces down to 3.5 times more candidates selected by MB25 than the diffusion model (Table 5).
The left panel of Fig. 17, shows the AGN candidates selected by the diffusion, but predicted to have low AGN fraction by the model in MB25. Whereas the right panel shows the AGN fraction of agreed sources between the two methods. Fig. 26 shows where the MB25 selections are located on the inpainting metrics. As shown in Table 4, many of the MB25 sources are not located above the diffusion threshold. Of the sources that lie above the threshold, the majority show an . This implies that the selection is more sensitive to fainter AGN.
| Matched sources | MB25 AGN | MB25 Non-AGN |
| Diffusion selected | 2 263 | 1 331 |
| Diffusion non-selected | 19 758 | 402 429 |
| Method | Diffusion | MB25 |
| Total AGN Candidates | 16 053 | 57 874 |
Comparing the predictions for each of the Euclid-based selections in Fig. 18, we can see that our diffusion model covers areas of the colour space that the photometry-based selections do not. Being able to select both bluer galaxies, as captured by the photometry selections (MZ25; Euclid Collaboration: Bisigello et al. 2024), and the much redder sources captured by MB25, shows the versatility of the method, highlighting its effectiveness across different galaxy populations.
7 Summary and conclusions
In this paper, we have used a diffusion-based inpainting model to identify AGN and QSOs using VIS images from the Q1 data (Euclid Quick Release Q1 2025) from the Euclid telescope. Our approach focuses on inpainting the brightest pixels within the centre of a galaxy, and employs a novel thresholding approach based on the reconstruction errors to differentiate AGN and QSOs from the rest of the galaxy population. Using only VIS images, our method can generate a large, reliable sample of optically-selected AGN, without the requirement of AGN labels ahead of time. Utilising a standard training and inference pipeline with minimal modifications for astronomical data, our method demonstrates high recall and significant overlap with AGN selections obtained through more traditional methods, across various wavelengths including optical, near-infrared (NIR), mid-infrared, and Xray.
7.1 Considerations on training and inference costs
Our DDPM inpainting model required approximately 50 hours of training using a single Nvidia A6000 GPU, which aligns with expectations for models of similar scale and architecture. However, the inference pipeline is heavily constrained by memory rather than computation, as demonstrated by its near-linear scaling – the inpainting of 1024 and 8192 images took roughly the same amount of time when sufficient GPU memory was available. This memory constraint is due to the need to store and iteratively manipulate large intermediate tensors during inference. Whilst training, although more computationally expensive, has a more structured process where gradients are computed and accumulated in a predictable manner.
With an expanded compute resource (4 Nvidia H100 GPUs), we are able to inpaint just over 2.5 million images per day. As we have shown in Sects. 4 and 2.3, our model architecture and hyperparameters remain largely unmodified from their standard configurations. Optimisations tailored explicitly for astronomical-based data, especially in regards to the noise scheduler, will significantly reduce the number of iterations and inference time required to process each image. For example, Fig. 7 shows an opportunity for a two- to four-fold speed-up in timesteps when correcting for the high number of pure Gaussian noise images alone.
The results from Sect. 6.6 show that although our predictions have a significant overlap with many of the selections presented, it also highlights that the ROC classifier produces similar overlap performance. Given this boundary is formed by a simple calculation on the data, rather than requiring a complex model, it may present itself as the preferred method. However, it must be reiterated that the current pipeline has had minimum optimisations for the data and the classification task. Therefore, the ROC classifier serves as a fundamental baseline for comparison in our study. Future enhancements to the model are expected to not only improve the accuracy of identifying AGN candidates but also refine our reconstruction error maps for better analysis of prospective AGN components. The other classifiers do not offer the same potential for improvement, making the diffusion method the most promising moving forward.
Our key contributions and conclusions include the following
Compared to traditional colour-selection criteria, the approach is attaining high coverage in AGN candidates while achieving near-perfect separation of galaxies not featuring an AGN component. Reliably separating these different sources allows for creation of subsets with high purity which are vital for the cosmological objectives of the Euclid mission.
By demonstrating that VIS images alone can generate reliable and extensive samples of AGN and QSOs, our method showcases the impressive capabilities of Euclid’s imaging. Given the future of the mission will cover , the ability for AGN identification using only VIS allows for the possibility of some science aims to be met without the immediate need for integration of multiple data sources. This can significantly reduce time and complexity for processing and analysing the data.
The application of diffusion models with minimal adaptations for astronomical imaging highlights the potential for the use of more advanced machine-learning tools and pipelines for future Euclid data, as well as other scientific data sets. The ability to leverage models that can process large volumes of data efficiently, without costly and extensive specialised tuning, will be vital as scientific data sets continue to increase in size and complexity.
Although our model performs well without significant optimisation for our use case, we provide insights into the behaviour of the model given the challenges presented with astronomical imaging. We show how the assumptions of machine-learning literature can result in non-optimal training, especially when it comes to the noise scheduler and training losses. By highlighting these inefficiencies, we can see where progress can be made in both astronomy and machine-learning domains.
Because this is the first use of these pipelines for such an application, we have routinely provided guidance on improvements and next steps to refine this work. Enhancing the model’s precision, improving the efficiency of both training and inference, as well as expanding its capabilities, will only increase its impact.
In conclusion, our results demonstrate the feasibility and effectiveness of using diffusion-based inpainting for AGN and QSO identification. The ability to use VIS images exclusively for selecting these candidates, provides a potential strategic flexibility to the Euclid mission whilst aiding in reducing the reliance on complex multispectral data integration. By proving the utility of applying our pipelines to the task of AGN identification, we believe it can be expanded and can significantly contribute to a broader range of astronomical applications, such as identifying gravitational lensing or finding transient objects.
Acknowledgements.
[…]
- Adam et al. (2022) Adam, A., Coogan, A., Malkin, N., et al. 2022, arXiv preprint arXiv:2211.03812
- Adam et al. (2023) Adam, A., Stone, C., Bottrell, C., et al. 2023, arXiv preprint arXiv:2311.18002
- Assef et al. (2018) Assef, R. J., Stern, D., Noirot, G., et al. 2018, ApJS, 234, 23
- Davis et al. (2018) Davis, B. L., Graham, A. W., & Cameron, E. 2018, ApJ, 869, 113
- Davis et al. (2019) Davis, B. L., Graham, A. W., & Cameron, E. 2019, ApJ, 873, 85
- DESI Collaboration et al. (2024) DESI Collaboration, Adame, A. G., Aguilar, J., et al. 2024, AJ, 168, 58
- Dhariwal & Nichol (2021) Dhariwal, P. & Nichol, A. 2021, Advances in neural information processing systems, 34, 8780
- Dressler (1989) Dressler, A. 1989, in IAU Symposium, Vol. 134, Active Galactic Nuclei, ed. D. E. Osterbrock & J. S. Miller, 217
- Euclid Collaboration: Bisigello et al. (2024) Euclid Collaboration: Bisigello, L., Massimo, M., Tortora, C., et al. 2024, A&A, 691, A1
- Euclid Collaboration: Cropper et al. (2024) Euclid Collaboration: Cropper, M., Al Bahlawan, A., Amiaux, J., et al. 2024, A&A, accepted, arXiv:2405.13492
- Euclid Collaboration: Margalef-Bentabol et al. (2025) Euclid Collaboration: Margalef-Bentabol, B., Wang, L., La Marca, A., et al. 2025, A&A, submitted
- Euclid Collaboration: Matamoro Zatarain et al. (2025) Euclid Collaboration: Matamoro Zatarain, T., Fotopoulou, S., Ricci, F., et al. 2025, A&A, submitted
- Euclid Collaboration: Mellier et al. (2024) Euclid Collaboration: Mellier, Y., Abdurro’uf, Acevedo Barroso, J., et al. 2024, A&A, accepted, arXiv:2405.13491
- Euclid Collaboration: Romelli et al. (2025) Euclid Collaboration: Romelli, E., Kümmel, M., Dole, H., et al. 2025, A&A, submitted
- Euclid Collaboration: Roster et al. (2025) Euclid Collaboration: Roster, W., Salvato, M., Buchner, J., et al. 2025, A&A, submitted
- Euclid Collaboration: Walmsley et al. (2025) Euclid Collaboration: Walmsley, M., Huertas-Company, M., Quilley, L., et al. 2025, A&A, submitted
- Euclid Quick Release Q1 (2025) Euclid Quick Release Q1. 2025, https://doi.org/10.57780/esa-2853f3b
- Ferrarese & Merritt (2000) Ferrarese, L. & Merritt, D. 2000, ApJ, 539, L9
- Gaia Collaboration: Bailer-Jones et al. (2023) Gaia Collaboration: Bailer-Jones, C. A. L., Teyssier, D., Delchambre, L., et al. 2023, A&A, 674, A41
- Gaia Collaboration: Vallenari et al. (2023) Gaia Collaboration: Vallenari, A., Brown, A. G. A., Prusti, T., et al. 2023, A&A, 674, A1
- Goodfellow et al. (2014) Goodfellow, I., Pouget-Abadie, J., Mirza, M., et al. 2014, Advances in neural information processing systems, 27
- Goodfellow et al. (2020) Goodfellow, I., Pouget-Abadie, J., Mirza, M., et al. 2020, Communications of the ACM, 63, 139
- Harrison & Ramos Almeida (2024) Harrison, C. M. & Ramos Almeida, C. 2024, Galaxies, 12, 17
- Ho et al. (2020) Ho, J., Jain, A., & Abbeel, P. 2020, Advances in neural information processing systems, 33, 6840
- Karras et al. (2020) Karras, T., Laine, S., Aittala, M., et al. 2020, in IEEE/CVF conference on computer vision and pattern recognition (IEEE), 8110–8119
- Kingma (2013) Kingma, D. P. 2013, arXiv preprint arXiv:1312.6114
- Kormendy & Richstone (1995) Kormendy, J. & Richstone, D. 1995, ARA&A, 33, 581
- Kullback & Leibler (1951) Kullback, S. & Leibler, R. A. 1951, The annals of mathematical statistics, 22, 79
- Lugmayr et al. (2022) Lugmayr, A., Danelljan, M., Romero, A., et al. 2022, in IEEE/CVF conference on computer vision and pattern recognition (IEEE), 11461–11471
- Lupton et al. (2004) Lupton, R., Blanton, M. R., Fekete, G., et al. 2004, PASP, 116, 133
- Magorrian et al. (1998) Magorrian, J., Tremaine, S., Richstone, D., et al. 1998, AJ, 115, 2285
- Nichol & Dhariwal (2021) Nichol, A. Q. & Dhariwal, P. 2021, in Proceedings of Machine Learning Research, Vol. 139, Proceedings of the 38th International Conference on Machine Learning, ed. M. Meila & T. Zhang (PMLR), 8162–8171
- Pathak et al. (2016) Pathak, D., Krahenbuhl, P., Donahue, J., Darrell, T., & Efros, A. A. 2016, in IEEE conference on computer vision and pattern recognition (IEEE), 2536–2544
- Radford et al. (2021) Radford, A., Kim, J. W., Hallacy, C., et al. 2021, in Proceedings of Machine Learning Research, Vol. 139, Proceedings of the 38th International Conference on Machine Learning, ed. M. Meila & T. Zhang (PMLR), 8748–8763
- Ramesh et al. (2022) Ramesh, A., Dhariwal, P., Nichol, A., Chu, C., & Chen, M. 2022, arXiv e-prints, arXiv:2204.06125
- Remy et al. (2023) Remy, B., Lanusse, F., Jeffrey, N., et al. 2023, A&A, 672, A51
- Saharia et al. (2022) Saharia, C., Chan, W., Saxena, S., et al. 2022, in Advances in Neural Information Processing Systems, ed. S. Koyejo, S. Mohamed, A. Agarwal, D. Belgrave, K. Cho, & A. Oh, Vol. 35 (Curran Associates, Inc.), 36479–36494
- Sahu et al. (2019) Sahu, N., Graham, A. W., & Davis, B. L. 2019, ApJ, 876, 155
- Sampson et al. (2024) Sampson, M. L., Melchior, P., Ward, C., & Birmingham, S. 2024, Astronomy and Computing, 49, 100875
- Sargsyan et al. (2023) Sargsyan, A., Navasardyan, S., Xu, X., & Shi, H. 2023, in IEEE/CVF International Conference on Computer Vision (IEEE), 7335–7345
- Smith et al. (2022) Smith, M. J., Geach, J. E., Jackson, R. A., et al. 2022, MNRAS, 511, 1808
- Spagnoletti et al. (2024) Spagnoletti, A., Boucaud, A., Huertas-Company, M., Kabalan, W., & Biswas, B. 2024, arXiv preprint arXiv:2411.19158
- Stone & Courteau (2019) Stone, C. & Courteau, S. 2019, ApJ, 882, 6
- Stone et al. (2021) Stone, C., Courteau, S., & Arora, N. 2021, ApJ, 912, 41
- Storey-Fisher et al. (2024) Storey-Fisher, K., Hogg, D. W., Rix, H.-W., et al. 2024, ApJ, 964, 69
- Thanh-Tung & Tran (2020) Thanh-Tung, H. & Tran, T. 2020, in 2020 international joint conference on neural networks (ijcnn), 1–10
- Wang et al. (2018) Wang, Y., Tao, X., Qi, X., Shen, X., & Jia, J. 2018, Advances in neural information processing systems, 31
- Yeh et al. (2017) Yeh, R. A., Chen, C., Yian Lim, T., et al. 2017, in IEEE conference on computer vision and pattern recognition (IEEE), 5485–5493
- Yu et al. (2018) Yu, J., Lin, Z., Yang, J., et al. 2018, in IEEE conference on computer vision and pattern recognition (IEEE), 5505–5514
- Zhang et al. (2023) Zhang, L., Rao, A., & Agrawala, M. 2023, in IEEE/CVF International Conference on Computer Vision (IEEE), 3836–3847
Appendix A Reconstruction metrics
Figures Fig. 19 show the differences in the rate of change measurement between input images and the inpainting output.
Appendix B Extreme examples
We present output examples from the inpainting pipeline where pixel values deviate significantly from the expected input pixels. Figure 20 highlights much fainter pixels, while Figure 21 shows much brighter ones.
Appendix C Morphology examples
We show output examples from various morphology types to showcase how the diffusion model adapts to different galaxy shapes. Edge-on galaxies, spirals and mergers are shown in Fig. 22,Fig. 23 and Fig. 24, respectively.
Appendix D Additional results
We present collated results for each of the star, galaxy, AGN and QSO selections in Fig. 25. In Fig. 26, the MSE and max ratio scores of all the MB25 selections as shown, indicating our diffusion model is more sensitive to selecting brighter sources.

### v2

@@
 2018, 2019 and references therein), it has been believed that most – if not all – galaxies with a spheroidal (or pseudo-spheroidal) component contain a SMBH at their centre.
-If galaxy’s emission shows evidence for its central SMBH being actively fuelled, it will emit non-thermal radiation leading it to be classified as an active galactic nucleus (AGN).
+If a galaxy’s emission shows evidence for its central SMBH being actively fuelled, it will emit non-thermal radiation, leading it to be classified as an active galactic nucleus (AGN).
 Depending upon the luminosity of that emission and its contrast with the stellar emission from the surrounding galaxy, it could have a classification ranging from a Seyfert galaxy (for the weakest emission) through to a quasi-stellar object or QSO (for the most luminous).
@@
 The growth and mass of a central SMBH is intimately related to the evolution of its host galaxy, not least through the action of energy, momentum and radiative feedback from outflows originating from the area around the SMBH on the properties of the stellar populations and ISM (see Harrison & Ramos Almeida 2024 for a recent review).
-Consequently, identifying and characterising SMBHs, particularly during periods of active feedback is central to the study of SMBH and galaxy evolution.
-A more precise classification of a source as an AGN is possible by using a much wider range of multi-wavelength data.
+Consequently, identifying and characterising SMBHs, particularly during periods of active feedback, is central to the study of SMBH and galaxy evolution.
+A more precise classification of a source as an AGN is possible when a much wider range of multi-wavelength data is available.
 This can include spectroscopy and considering the much broader spectral energy distribution of the galaxy and/or the candidate AGN.
@@
 The technique can then be subsequently refined by feeding back the result of using other multi-wavelength data, allowing us to confirm or reject candidates.
-With sufficient refinement, this further assessment step subsequently becomes either unnecessary or at the very least an efficient stage in compiling reliable samples.
-This then opens the door to production of unprecedentedly large and reliable samples of optically-selected AGN.
-In this work, we present a technique that can achieve the above goal.
-In the following we describe the use of diffusion models for generating samples of AGN candidates using Euclid VIS band (Euclid Collaboration: Mellier et al.
-2024; Euclid Collaboration: Cropper et al.
-2024) imaging alone.
-Diffusion-based methods are machine-learning models used to generate or reconstruct data by evolving a random distribution of pixels into a structured output over several steps.
+With sufficient refinement, this further assessment step eventually becomes either unnecessary, or at the very least, an efficient stage in compiling reliable samples.
+This then opens the door to the production of unprecedentedly large and reliable samples of optically-selected AGN.
+This work presents a technique that can achieve the above goal.
+In the following, we describe the use of diffusion models for generating samples of AGN candidates using Euclid VIS band (Euclid Collaboration: Mellier et al.
+2025; Euclid Collaboration: Cropper et al.
+2025) imaging alone.
+Diffusion-based methods are machine-learning models that generate or reconstruct data by evolving a random distribution of pixels into a structured output over several steps.
 This technique outperforms traditional methods like median pixel substitution or interpolation, which often struggle to recreate complex structures within images.
-By using these models for ‘inpainting’, where we adaptively recreate parts of the image, we can measure the discrepancies between the original and generated pixels.
+Using these models for ‘inpainting’, where we adaptively recreate parts of the image, we can measure the discrepancies between the original and generated pixels.
 Examining these errors allows us to separate AGN, QSOs, and other sources from the broader galaxy population.
-Briefly, the technique takes an image of a galaxy that may or may not have AGN emission at its centre and excludes those central pixels that may be influenced by AGN emission.
-The model then predicts (‘inpaints’) the underlying stellar light profile in those pixels from a comparison of the rest of the galaxy image to a model derived from a very large sample of similar galaxy images.
-Any significant difference between the prediction and true image in those central regions indicates the potential presence of AGN emission.
-Because AGN are comparatively rare, the sample of galaxy images used in the generation of the model does not need to be filtered to reject AGN, the method will be inherently biased to the ‘typical’ behaviour of galaxy light distributions and so naturally down-weights or excludes predicting central regions with AGN emission.
-While this paper details the basic technique applied to VIS imaging data, as this is a machine-learning technique, future enhancements would allow incorporation of the results of a detailed assessment step, involving other data sets, in order to further improve the accuracy and reliability of the AGN candidate samples produced directly by the technique.
-Such steps could include multi-wavelength cross-matching or spectroscopic validation, from which we can ultimately reduce the need for further time-expensive refinement steps.
-Additionally, the technique as presented here uses a standard treatment of diffusion models which in principle can be significantly further optimised for the specific characteristics of the Euclid data set, thereby improving further the efficiency of the processing of the Euclid data.
+Our method identifies potential active galactic nuclei (AGN) by treating the task as an unsupervised anomaly detection problem.
+It begins with an image of a galaxy, potentially containing AGN emission at its center, and masks out the central pixels that might be influenced by such emission.
+A diffusion model, trained on a large, unlabeled data set of galaxies – most of which are normal, non-active – then predicts or ‘inpaints’ the masked region based on the learned prior distribution of typical galaxy morphologies.
+Because AGN are rare, there is no requirement to remove them from the training sample; the model is inherently biased toward reconstructing the central region as if it belonged to a typical, non-active galaxy.
+As a result, any significant discrepancy between the inpainted and actual central region suggests the presence of an AGN.
+This reconstruction error thus serves as a powerful anomaly score for identifying candidate AGN.
+While this paper details the fundamental technique applied to VIS imaging data, as this approach utilises machine learning, future improvements could incorporate findings from a detailed assessment that involves additional data sets.
+These could include multi-wavelength cross-matching or spectroscopic validation, from which we can ultimately reduce the need for further time-expensive refinement steps.
+Such enhancements would help to further improve the accuracy and reliability of the AGN candidate samples generated by the technique.
+Additionally, the technique as presented here uses a standard treatment of diffusion models, which in principle can be significantly further optimised for the specific characteristics of the Euclid data set, thereby improving further the efficiency of the processing of the Euclid data.
 Such work would be the subject of future papers.
@@
 2.1 Generative models
-Generative models such as generative adversarial networks (GANs, Goodfellow et al.
-2014, 2020) or variational autoencoders (VAEs, Kingma 2013) are trained with the aim to learn the underlying distribution of a data set to generate novel examples.
-Using a pair of competing networks, GANs ensure that once trained the generated images are of high quality, leading to the creation of many artificially generated data sets that are indistinguishable from the original data set (Karras et al.
-2020).
-However, the adversarial training loop is often difficult to train, often leading to mode collapse, meaning the generator focuses on only a small subset of the total search space leading to generating samples of similar instances (Thanh-Tung & Tran 2020).
-On the other hand, VAEs use an encoder network to compress the data into a latent space of fewer dimensions.
-The network then uses the latent space to recreate the input through an often symmetric decoder network.
-VAEs provide better coverage over the search space, improving stability during training as well as increased variety in generated outputs.
-However, the drawback for VAEs becomes apparent when handling more complex, high dimensional data.
-The use of latent spaces does not guarantee a lossless compression of information, often leading to accurate but blurry outputs.
+Generative models are a class of machine learning systems designed to learn an underlying data distribution, , from a finite set of independent and identically distributed (i.i.d.) samples .
+The primary goal of these models is to develop a tractable method for generating new, synthetic samples that are statistically indistinguishable from the training data.
+Over the years, several distinct families of generative models have emerged, each with unique architectural and training paradigms.
+A key distinction lies in their optimisation objectives, which often seek to minimise a measure of divergence between the true data distribution and the model’s learned distribution .
+The choice of divergence has significant implications for model behavior.
+Prominent examples include Variational Autoencoders (VAEs, Kingma 2013), which learn a latent variable model of the data, and Generative Adversarial Networks (GANs, Goodfellow et al.
+2014, 2020), which involve a contest between a generator and a discriminator.
+GANs implicitly optimise an objective related to the reverse Kullback-Leibler (KL) divergence, , which exhibits a ‘mode-seeking’ property.
+This encourages the generation of high-fidelity samples but can lead to mode collapse, where the model learns only a subset of the data distribution (Thanh-Tung & Tran 2020).
+In contrast, VAEs and Diffusion Models optimise objectives related to the forward KL divergence, , which is ‘mass-covering’.
+This property encourages the model to account for all modes present in the data, leading to greater sample diversity and more comprehensive coverage of the data distribution.
 2.2 Diffusion models
-Diffusion-based models provide an alternative process for image generation that allows for both good coverage, as well as the ability to represent complex data distributions (Dhariwal & Nichol 2021).
-For both GANs and VAEs, their significant advantage with respect to diffusion models is that they offer faster sample generation and lower computational costs.
-Whereas diffusion models trade off speed for superior sample quality, mode coverage, and flexibility in controllable generation making them our preferred choice in the task of AGN identification.
+Diffusion Models (Sohl-Dickstein et al.
+2015; Ho et al.
+2020), the focus of this work, systematically corrupt data with noise and then learn to reverse the process.
+They have been shown to achieve superior sample quality and mode coverage compared to other classes of models (Dhariwal & Nichol 2021), making them a compelling choice for modelling the complex and diverse morphologies of galaxies observed by Euclid.
+For both GANs and VAEs, their significant advantage compared to diffusion models is that they offer faster sample generation and lower computational costs.
+Conversely, diffusion models trade off speed for superior sample quality, mode coverage, and flexibility in controllable generation, making them our preferred choice in the task of AGN identification.
 This paper makes use of one such implementation, denoising diffusion probabilistic models (DDPMs, Ho et al.
 2020).
-Diffusion works on the following premise: by continually adding a small amount of noise to an image, eventually one is left with an image of complete noise without any remnants of the original input.
+Diffusion works on the following premise: by continually adding a small amount of noise to an image, eventually, one is left with an image of complete noise without any remnants of the original input.
 If the noise is generated in a stochastic but consistent way, such as from a Gaussian distribution, we can use a network to learn the dynamics of the noise and perform a reverse process.
@@
 The term defines the variance of the Gaussian noise added at each timestep.
-By ensuring a gradual increase in noise, the model is more easily able to learn the transition from the data distribution to the noise distribution.
+By ensuring a gradual increase in noise, the model can more easily learn the transition from the data distribution to the noise distribution.
 With a large enough and an adjusted time-dependent variance , will approach a Gaussian distribution that is isotropic, meaning uniform in all directions.
@@
 1.
+Background noise versus diffusion noise
+It is important to clarify the distinctions between the definition and use of ‘noise’ in diffusion models compared to the traditional understanding of noise in astronomy-based data.
+In the context of astronomical imaging, noise refers to the random fluctuations inherent in observational data, often caused by e.g.
+background emission, electronics read-out noise, and atmospheric disturbances.
+This ‘background noise’ reduces the signal-to-noise ratio (S/N) of the image, affecting the ability to detect subtle details in any analysis.
+In contrast, the diffusion process uses an intentional integration of Gaussian noise to facilitate the training of the generative model.
+The pixels of image , follow the distribution: , rather than following the distribution of the background noise.
+The goal of the reverse diffusion process is to remove the artificial, high-variance Gaussian noise that was intentionally added during the forward process, thereby transforming a sample from a simple prior distribution () into a realistic galaxy image.
+A realistic galaxy image, as defined by our training data set, inherently includes the source’s light profile superimposed on the characteristic background noise of the Euclid instrument.
+Therefore, the model learns to generate images that replicate this entire distribution, including the statistical properties of the background noise.
+The diffusion pipeline does not aim to produce a ‘denoised’ image in the traditional astronomical sense (i.e., a background-subtracted image).
+Use in Astronomy Applications
 Smith et al.
@@
 2021), a collection of large, well resolved objects that feature significant internal structure.
-They provide analysis of the quality of their generated images, showcasing the diffusion model’s ability to create not only visually realistic images, but also how they feature similar physical property distributions such as half-light radius and flux-space colour values.
+They provide analysis of the quality of their generated images, showcasing the diffusion model’s ability to create visually realistic images which feature similar physical property distributions such as half-light radius and flux-space colour values.
 They also briefly explore the use of inpainting with their model to remove satellite trails.
@@
 It is for this reason that score-based models can be classed as the more general framework for diffusion-based generative models.
-The use of score-based generative models within astronomical applications include galaxy image deconvolution (Adam et al.
+The use of score-based generative models within astronomical applications includes galaxy image deconvolution (Adam et al.
 2023; Spagnoletti et al.
@@
 2024).
+2.3 Training Objective
+For training, we use 6464 pixel VIS cutouts for each of the selected sources that will be discussed in Sect.
+4.1.
+We use the training pipeline described in Dhariwal & Nichol (2021)111Available at https://github.com/openai/guided-diffusion and adopt a hybrid loss function that combines a simple mean squared error (MSE) term with the full variational lower bound (VLB).
+The MSE objective, proposed by Ho et al. (2020), is an MSE loss between the true and predicted Gaussian noise,
+where is an original image, is a uniformly sampled timestep, and is the true noise.
+For our implementation is described below, where the reader is directed to Nichol & Dhariwal (2021) for specifics on its derivation
+While this objective is effective for training the mean of the reverse process transition, , it provides no learning signal for its variance, .
+To address this, Nichol & Dhariwal (2021) proposed learning this variance by incorporating the full VLB of the log-likelihood into the training objective.
+The VLB is given by
+Since explicitly depends on the parameters of , it can be used to train them. We adopt the hybrid objective
+We use a small weighting constant () to prevent the term from overpowering the term, which remains the primary driver of sample quality.
+This hybrid loss allows the model to learn an optimal reverse process variance, which can improve model likelihoods and enable more efficient sampling.
+A key challenge with astronomical data is its high dynamic range.
+A standard MSE-based loss is not scale-invariant and will be dominated by the few brightest pixels or sources in a training batch.
+This can lead to poor performance on the much larger population of fainter objects.
+To address this, we introduce a normalised version of the MSE loss term.
+For each image , we divide its error by its respective maximum pixel value, denoted as .
+This normalisation ensures that the relative error is penalised, allowing for more equitable training across both bright and faint sources.
+The final loss used for our primary model is thus
+Throughout the remainder of the paper, we refer to this variant of the loss as the normalised loss.
+The training curves of this scale-invariant loss compared with the original loss used in Nichol & Dhariwal (2021) is shown in Fig.
+18.
+By incorporating this normalised loss, we enable more equitable training across the full range of galaxy brightnesses, rather than being dominated by the brightest objects which occurs when using the default loss.
 2.4 Noise scheduling
-The larger the number of total timesteps , the more refined the model can become at accurately generating realistic images, since less complex noise dynamics are required to be learnt for each timestep .
-This does, however, require the model to learn the dynamics of more timesteps, leading to extra overhead in both training and inference, since more steps must be performed to traverse from pure noise to realistic images.
+The total number of timesteps, , is a critical hyperparameter in the diffusion process.
+A larger corresponds to smaller, more refined steps in both the forward (noising) and reverse (denoising) processes.
+While this can, in principle, allow the model to learn a more accurate representation of the data distribution, it comes with a significant trade-off.
+The computational cost of inference (i.e., generating a sample) is directly proportional to , as the reverse process must be iterated times.
+The cost of a single training step, however, is not directly dependent on , because at each step a random timestep is sampled and the loss is evaluated only for that level of noise.
 In the original DDPM paper, Ho et al.
 (2020) made use of a linear schedule for the amount of noise added to the image at each .
-Although the produced images were competing with state-of-the-art methods, Nichol & Dhariwal (2021) found that a linear schedule was not well suited to low resolution images and proposed using a cosine-beta noise schedule as a replacement.
-By stretching the noise levels so that less noise is added in early iterations and more in the later, the process overcame the issue of images became too noisy too quickly.
+Although the produced images competed with state-of-the-art methods, Nichol & Dhariwal (2021) found that a linear schedule was not well suited to low-resolution images and proposed using a cosine-beta noise schedule as a replacement.
+By stretching the noise levels to add less noise in early iterations and more in the latter, the process overcame the issue of images becoming too noisy too quickly.
 Figure 2 shows how the scheduler impacts the parameter used in Eq.
 2.
-Equation 3 shows how is calculated, where the reader is directed to Nichol & Dhariwal (2021) for specifics on its derivation:
-where
-All models in this paper make use of the cosine-beta scheduler. The impact of this scheduler on the high dynamic range nature of astronomy data is explored in detail in Sect. 5.
+Equation 10 shows how is calculated, where the reader is directed to Nichol & Dhariwal (2021) for specifics on its derivation
+where is shown in Eq. 4.
+All models in this paper make use of the cosine-beta scheduler. The impact of this scheduler on the high dynamic range nature of astronomy data is explored in detail in Sect. 6.1.
 2.5 Conditioning the model
 Using the pipeline of DDPMs allows us to impose conditioning on our input space.
-Throughout training, the model is already conditioned on .
-Due to the cosine scheduler discussed previously, the model will alter an image very differently at a timestep close to , where the image is still very noisy, than when the image is close to timestep and near completion.
-However, we are not limited to only conditioning by timestep, often multiple parameters are provided alongside each image.
-Typically, this is in respect to ’class-conditioning’, where users can tune the expected output of the diffusion model to a specific class or category (Radford et al.
+Throughout training, the model is already conditioned on the timestep .
+This conditioning is crucial, as it informs the neural network of the noise level present in the input image .
+The network learns a single function, , that is trained to predict the noise component from a noised image at any given timestep .
+The characteristics of the noise to be removed differ substantially between a timestep close to (where the image is almost pure noise) and a timestep close to (where the image is almost fully reconstructed).
+Providing as an input allows the single network to adapt its predictions accordingly.
+However, we are not limited to only conditioning by timestep; often, multiple parameters are provided alongside each image.
+Typically, this is to allow ‘class-conditioning’, where users can tune the expected output of the diffusion model to a specific class or category (Radford et al.
 2021; Zhang et al.
 2023).
-By providing an embedding space during training, the model learns not only the dynamics of the noise but the inherent differences of noise and non-noise pixel values for each of the representations.
+By providing an embedding space during training, the model can learn not only the dynamics of the noise but also the inherent differences between noise and non-noise pixel values for each of the class representations.
 This conditioning provides small nudges to the random walk, guiding the denoising to the respective population of the class.
@@
 This type of conditioning allows the user to direct the image generation to a particular class or style of image.
-However, with the possibility of complex prompts and classes, it may be difficult to know exactly what the output image will look like.
-Furthermore, with diffusion’s ability to cover large areas of the search space, a repeat sample of the same prompt is likely to produce an image very different from the first, even if the generated image belongs to the requested class.
+However, with diffusion’s ability to cover large areas of the search space, a repeat sample of the same prompt will likely produce an image very different from the first, even if the generated image belongs to the requested class.
+Even with the possibility of detailed prompts and refined classes, which can guide the generation towards a narrower or more specific region of the learned space, it may be difficult to know exactly what the output image will look like.
 This is due to the one-to-many relationship between classes/prompts and their subsequent generated outputs.
-In a situation where the user requires consistency in part of the image, a more direct manipulation of the produced pixels is required.
-A process known as ’inpainting’ (Pathak et al.
+However, for our task, we require a different form of conditioning.
+Instead of generating an entire image based on a class label, we need to reconstruct a specific, masked region of an existing image, conditioned on the surrounding, unmasked pixels.
+This task is known as inpainting (Pathak et al.
 2016; Yeh et al.
@@
 2018; Wang et al.
-2018), forces conditioning of specific pixels within the image, allowing the model to fix certain parts of the image whilst generating others – all whilst remaining coherent and consistent across the image as a whole.
-Allowing the user to selectively remove and replace certain aspects of images has gained a lot of popularity in the last few years due to its mainstream adoption by nearly all smartphone brands (Sargsyan et al.
-2023).
+2018).
+To achieve this, we do not train a new, specialised conditional model.
+Instead, we employ a ‘plug-and-play’ (PnP) approach, where conditioning is introduced during the inference phase of an unconditionally trained generative model (Nguyen et al.
+2017; Graikos et al.
+2022).
+Our diffusion model, trained on a vast sample of galaxies, serves as a powerful, non-parametric prior for galaxy morphology.
+The PnP approach allows us to apply this prior to the inpainting task with great flexibility, as the model does not need to be retrained for different images or mask geometries.
+This is particularly advantageous for large-scale surveys like Euclid.
+Recent works in astrophysics have successfully employed similar PnP or posterior sampling strategies for complex inference tasks (Feng et al.
+2023; Dia et al.
+2025).
 2.6 Repainting
-This paper uses a specific implementation of inpainting using diffusion known as Repaint (Lugmayr et al.
-2022).
-Due to the model being conditioned on timestep , where it is expecting the whole image to be at a certain noise level, we can not simply take an image and a corresponding mask and replace the masked pixels with noise.
-The model itself does not know that a mask has been applied nor does it know that the inverse mask should remain unaffected.
-Instead, it will continue to remove noise across the whole image as usual, leading to the clean parts of the image becoming corrupted.
-It is also not possible to generate an entirely new image and copy over the masked pixels to the original image as there will be no continuity between the existing and new elements.
-To combine the parts of the image we want unchanged, and the parts we want to generate, both the forward pass and the backward pass of the diffusion process must be utilised.
-This pipeline needs to generate an image that incorporates the mask and the fixed pixels for each timestep , as shown in Fig.
+For this work, we use the Repaint algorithm (Lugmayr et al.
+2022), a PnP method specifically designed for diffusion models.
+The core challenge in diffusion-based inpainting is to ensure semantic and structural consistency between the known (unmasked) and unknown (masked) regions.
+A naive approach of simply replacing the masked region with noise and then running the reverse process would fail, as the denoising network would alter the known, unmasked pixels, corrupting the conditioning information.
+Repaint overcomes this by iteratively enforcing the data constraint at each step of the reverse diffusion process.
+Given an image and a binary mask (where indicates the region to be inpainted and indicates the known region), the algorithm proceeds as follows.
+At each timestep of the reverse process, starting from a noisy sample :
 1.
-Input
-The main input of each timestep of the repaint pipeline is the denoised image, , from the previous timestep.
-To aid in the understanding of the pipeline, this image is referred to as for this discussion.
-At the first iteration when , the input is simply an image of pure Gaussian noise.
-Fixed pixels
-To ensure that the pixels we want unchanged are still present at the end of the diffusion process, we must interject these fixed pixels during each timestep.
-However, as the model is expecting all pixels to be at a specific noise level, as described in Eq.
-2, we must utilise the same scheduler used in training to add the correct amount of noise for the current timestep .
-Due to the stochasticity present in adding Gaussian noise to the image, the pipeline allows us to create different versions of the input with the required noise.
-By averaging these samples, the important details of the image are more consistently preserved, ensuring that once all iterations are complete, the fixed pixels look identical to the original input.
-We refer to the image formed from the average of the N samples as .
-Utilising the mask
-To combine the parts of the image we want to keep and the parts we want to generate, we utilise a mask , which is an array of equal size to our image, with pixel values of either zero or one.
-Pixels valued as one will be generated from scratch by the diffusion process, with pixels valued zero being replaced at each noise step with the fixed pixels in from the process described above.
-The two parts can then be combined into a coherent image and then passed back into the diffusion model to be further denoised and progress onto timestep ,
-By denoising the combination of pixels we want to keep with the generated pixels, we condition the diffusion process to ensure that in the final image, all pixels are coherent, being conditioned on the details found within the fixed pixels.
-Incorporating the Repaint method for AGN identification has the following benefits
-Diffusion methods have been proven their ability in many applications to accurately recreate complex data sets, allowing for realistic generation of galaxy morphology images.
-There are no constraints on the size and shape of masks used and, since Repainting is only applied at inference, no additional retraining is required.
-There exists a large imbalance of normal galaxies compared to AGN, allowing for the model to learn a bias.
-This can be exploited to create an outlier detection-based classifier without explicitly requiring labels during training.
+Sample the known region: The original image, , is noised to the appropriate level for timestep using the forward process , resulting in a sample of the known part, .
+2.
+Sample the unknown region: The output image from the previous, slightly more noisy step, , is denoised for one step using the unconditional diffusion model to get a sample of the unknown part, .
+3.
+Combine and denoise: The known and unknown parts are combined: .
+This composite image is then passed to the next iteration of the repaint loop to produce the estimate for step .
+Once this image has been passed through one step of the learned reverse diffusion process , it will produce the slightly less noisy image .
+This procedure, illustrated in Fig.
+1, ensures that the generated region is coherently conditioned on the surrounding galaxy structure at every stage of the generation.
+The ‘denoise’ operation in this context refers to the application of one step of the learned reverse diffusion process, , which involves using the trained neural network to predict the noise in and deriving an estimate for .
+By denoising an image composed of both fixed and generated pixels – each matched to the appropriate noise level – we ensure that the generated region is properly conditioned on the information from the rest of the galaxy, resulting in a final image where all pixels are consistent and coherent.
+While alternative conditioning methods based on guiding the reverse process with a likelihood score function exist (Graikos et al.
+2022), we adopt the Repaint algorithm for its straightforward implementation and proven effectiveness in enforcing data consistency in a plug-and-play setting.
+Incorporating the Repaint method for AGN identification has the following benefits:
+Diffusion methods have proven their ability in many applications to accurately recreate complex data sets, allowing for the generation of realistic galaxy morphology images.
+A large imbalance of normal galaxies exists compared to AGN, allowing the model to learn a bias.
+Such a bias can be exploited to create an outlier detection-based classifier without explicitly requiring labels during training.
+There are no constraints on the size and shape of masks used, and since Repainting is only applied at inference, no additional retraining is required.
+3 Training and inference pipeline
+The diffusion and inpainting methods are two distinct pipelines, one for the initial training of the model and a second for the actual inpainting.
+Training the diffusion model uses the whole image and learns to predict what noise to remove at each timestep, allowing for generating novel examples from the trained distribution.
+3.1 Mask creation
+Due to the nature of the application, our aim is to mask out any suspected AGN without the need to rely on labels ahead of time.
+Each mask is positioned so that the brightest pixel in the centre of the source becomes the centre of the mask.
+Although each image is centred on the source, various interactions and dynamics allow suspected AGN to be offset from the centre.
+To allow for some deviation, a 99 pixel window around the centre of the image is used to detect the brightest pixel.
+This offset allows for the correct masking of less symmetric sources whilst limiting the likelihood of masking an adjacent source by mistake.
+The masks for each image are assigned automatically throughout the data-loading pipeline.
+Throughout this paper, we use a 55 pixel mask, unless explicitly stated otherwise, centred on the calculated brightest pixel.
+A 55 pixel mask, which covers 0505, allows us to effectively cover the AGN contribution whilst minimising loss of the surrounding galaxy structure.
+Given that the VIS PSF has a full width half maximum (FWHM) of 013 (Euclid Collaboration: Mellier et al.
+2025), the AGN’s light is spread over multiple pixels.
+Taking into account the redshift of the galaxy, AGN within nearby sources () will cover 4.8 pixels, ensuring that we can fully encapsulate the core without unnecessary overlap of the host morphology.
+By retaining as much of the host galaxy information as possible for inpainting, we also reduce the risk of introducing artificial features into the reconstructed image.
+3.2 Reconstruction rescaling
+Unlike some image processing pipelines that rescale image data to a limited range, such as when working with JPG images which are binned to the standard 0–255 pixel scale range and converted to floats between 0–1 (Walmsley et al.
+2020; Euclid Collaboration: Aussel et al.
+2024), our approach utilises the raw values direct from the sensor as input to our network.
+As comparisons of individual pixels are being made on such a small mask, the relative differences become much more impactful and would likely be lost if values were binned.
+One potential issue with this setup is that the libraries used for training are optimised for images within the standard pixel range.
+Although this did not prevent the model from effectively learning the distribution, it does require a rescaling of the generated output as the pixels are mostly between 1 and 1 to improve training stability.
+We experimented with simply normalising the input data so that all our images were in the range 0–1; however, due to the large dynamic range of the data, the faintest pixels ended up being .
+Given the previous discussions on how MSE does not penalise smaller errors effectively, the model could not accurately recreate sources, outputting a near-empty image of very low pixel values.
+To allow the inpainted region to be meaningfully compared (e.g., via MSE) to the original pixel values in that masked region, a post-inference rescaling must be applied.
+The vast majority of pixels in the image (the unmasked region, which constitutes 99.5% of a 6464 image when a 55 central mask is used) correspond to known pixel values from the original astronomical image.
+We leverage these known, unmasked pixels to perform a linear rescaling of the diffusion model’s initial output (which is in a normalised range) back to the original image’s dynamic range.
+This procedure ensures consistency in the unmasked regions and appropriately scales the newly inpainted pixels in the masked region.
+The following scaling can be applied instantly on output of the repaint procedure using each image’s respective values to allow the non-masked pixels of the input and output images to be consistent:
+where
+and
+By ensuring that the shared pixels are consistent, the inpainted pixels can also be correctly adjusted.
 4 Data
@@
 2025) in the Euclid Deep Fields North (EDF-N), South (EDF-S), and Fornax (EDF-F).
-In this work we use exclusively the VIS images (Euclid band), because we aim to identify AGN candidates on a single image.
+In this work, we use exclusively the VIS images (Euclid band) because we aim to identify AGN candidates on a single image.
 However, due to the varying distances and sizes of sources, an appropriate subset must be chosen to ensure the model is presented with a sufficiently broad selection of resolved galaxies.
@@
 the source must not have any defects or abnormalities as signified by the detection flags (det_quality_flag_euclid = 0 and flag_vis_euclid = 0);
-the source must have low probability of being spurious (spurious_prob_euclid 0.2);
+the source must have a low probability of being spurious (spurious_prob_euclid 0.2);
 the source must have a large segmentation area (segmentation_area_euclid200);
@@
 if the signal-to-noise criterion is not met, then the segmentation area of the source must be very large (segmentation_area_euclid 1200), as defined in Euclid Collaboration: Walmsley et al. (2025).
-After this sample selection we are left with 1 142 606 sources across the three Q1 fields.
-VIS image cutouts (6464 pixels) were created for each of the respective sources, where problematic images, such as those at the very edge of tiles where a full cutout could not be created, were removed.
+It is crucial to note that because AGN are intrinsically rare compared to the general galaxy population, the training set does not need to be pre-filtered to remove them.
+The diffusion model, in learning the data distribution, naturally develops a prior that is heavily biased towards the ‘typical’ light profiles of non-active galaxies.
+This inherent bias is the key mechanism that enables the subsequent outlier detection.
+After this sample selection, we are left with 1 142 606 sources across the three Q1 fields.
+VIS image cutouts (6464 pixels) were created for each source.
+Problematic images, such as those at the very edge of tiles where a full cutout could not be created, were removed.
 This provided us with images and catalogue data for 1 060 864 sources.
 We then apply a 80 : 20 train and test split.
-Even though the inpainting and its results are a separate pipeline to the diffusion training, to prevent any potential data leakage, all of the decision boundaries created are based solely on the training set.
+Even though the inpainting and its results are a separate pipeline to the diffusion training, all of the decision boundaries created are based solely on the training set to prevent any potential data leakage.
 When showing classifier performance, all results are from the unseen 20% of the data.
-This is to prevent any potential memorisation of the images from training impacting the results.
+This prevents any potential memorisation of the training images from impacting the results.
 However, in our analysis, no obvious differences have been found in the inpainting results between the two sets, showing that the model is not simply memorising the training set.
 To explore the impact point-like sources have on training and the inpainting process, a second smaller selection was created that combined all the selections above with the following:
-the source must have a low probability of being star. (phz_star_prob_phz_class 0.1); and
+The source must have a low probability of being a star. (phz_star_prob_phz_class 0.1); and
 the source must have a low probability of being point-like. (point_like_prob_euclid 0.2).
@@
 4.2 Comparisons with other classifications
-Although the training of the diffusion model does not make use of any labels, it is important to compare the distributions of metrics over different source selection criteria.
-This allows us to compare both the performance of detecting particular sources but also to determine what sources or factors may hinder the pipeline’s ability to classify objects sufficiently.
+Although our diffusion model is trained in an unsupervised manner without labels, it is essential to compare the distributions of metrics over different source selection criteria.
+This allows us to compare the performance of detecting particular sources and determine what sources or factors may hinder the pipeline’s ability to classify objects sufficiently.
 In particular, we use some of the sample definitions discussed in Euclid Collaboration: Matamoro Zatarain et al.
@@
 2018), which is referred to as R90 for the remainder of the paper.
-As is particularly true for AGN and QSO, it is impossible to create a selection criterion that will identify both with high completeness and high purity a single population of objects.
-Instead, we will use the aforementioned samples to assess contamination and purity of our method to the extent that is possible.
+The reader is directed to Euclid Collaboration: Matamoro Zatarain et al.
+(2025) for a comprehensive overview of each of the selections used throughout our paper.
+In this work, when referring to completeness against a specific reference sample (e.g., a catalog of selected AGN candidates), we mean the fraction of objects in that reference sample that are successfully identified by our method.
+This is due to the true astrophysical completeness being unknown and extremely difficult to measure from the observational data alone.
+Our context is equivalent to the standard machine learning metric of recall for the positive class, which is defined in Eq.
+13.
+When developing a method for initial candidate selection, where the goal is to flag objects for follow-up or multi-wavelength analysis, high recall is often prioritised.
+The aim is to minimise the loss of potentially interesting or rare sources (i.e., reduce false negatives), even if this comes at the cost of lower precision (i.e., including a larger fraction of contaminants or false positives in the selected sample).
+These contaminants can often be filtered out through subsequent analysis stages or by incorporating additional data.
+Our primary goal here is to demonstrate the method’s capability to recover a wide variety of known AGN and QSO types, making recall a key indicator of its potential utility in broad searches.
+As is particularly true for AGN and QSO, it is impossible to create a selection criterion that will identify a single population of objects with high completeness and high purity.
+Instead, we will use the aforementioned samples, many of which are only candidates themselves, to assess the contamination and purity of our method to the extent that is possible.
 4.3 Data expectations
-Although, by definition, a difference between a galaxy with and without a bright central component should be found, it is important to validate that this behaviour can be measured with the Euclid VIS images.
-To quantify this difference we analyse the relationship between the brightest central pixel and its surrounding pixels.
-By producing ratios comparing the central pixel to 1 and 2 pixels wide regions around the central pixel, we form a rate of change (ROC) metric over sources, as described by
+As a baseline for our morphology-based approach, it is first necessary to verify that the Euclid VIS images contain sufficient information to resolve the contrast between the compact central emission characteristic of an AGN and the broader light profiles of typical galaxies.
+We analyse the relationship between the brightest central pixel and its surrounding pixels to quantify a difference between normal galaxies and AGN.
+By producing ratios comparing the central pixel to 1 and 2-pixels-wide regions around the central pixel, we form a rate of change (ROC) metric over sources, as described by
 where
 and and are the mean values of the surrounding rings of 1 (total 8 pixels) and 2 pixels (total 16 pixels), respectively. is the pixel value of the brightest pixel within the centre 99 pixels.
-Pixels surrounding the brightest pixel, which share similar values to the maximum will produce a ROC value closer to zero, implying that the source does not feature a prominent active component.
-Sources featuring a single very bright pixel, surrounded by fainter pixels will have a ROC value closer to its maximum of 2.
+Pixels surrounding the brightest pixel, which share values similar to the maximum, will produce a ROC value closer to zero, implying that the source does not feature a prominent active component.
+Sources featuring a single very bright pixel surrounded by fainter pixels will have a ROC value closer to its maximum of 2.
 The produced distributions of sources, analysed according to both magnitude and various selection criteria (Sect.
-3.2), are shown in Fig.
+4.2), are shown in Fig.
 3.
-It is evident that sources with a high flux point-spread function (PSF) light profile show the highest ROC values, as expected.
-This effect is seen in the star sample and the QSO dominated selections (R90 WISE, Quaia, DESI-QSO, and Euclid-QSO).
+As expected, sources with a high flux point-spread function (PSF) light profile show the highest ROC values.
+This effect is seen in the star sample and the QSO-dominated selections (R90 WISE, Quaia, DESI-QSO, and Euclid-QSO).
 On the other hand, spectroscopically confirmed normal galaxies show very broad ROC distributions.
@@
 2004), the effect of using an transformation on the data is also investigated.
-The function, which is often to the benefit of visual inspections, reduces the differences between the brightest and faintest pixels.
-However, when attempting to differentiate galaxy and non-galaxy sources, the function becomes a hindrance, causing the distributions to more significantly overlap, as shown in Fig.
+The function is often used to benefit visual inspections by reducing the difference between the brightest and faintest pixels.
+In our approach, however, when attempting to differentiate galaxy and non-galaxy sources, the function becomes a hindrance, causing the distributions to more significantly overlap, as shown in Fig.
 4.
-To generate a measure for class overlap we measure the ROC for each image, taking the mean of each non-galaxy class.
-We use the percentage of our true galaxy sample that have a ROC value over the respective class mean as our overlap metric.
+To generate a measure for class overlap, we measure the ROC for each image, taking the mean of each non-galaxy class.
+We use the percentage of our true galaxy sample with a ROC value over the respective class mean as our overlap metric.
 The extent of the overlap is showcased in Table 1, where galaxy overlap increases by at least a factor of 2.5 and in the extreme case, nearly 50 times more galaxies overlap with stars.
-In the following sections of the paper we will discuss the set up, training and performance of the diffusion model that automatically learn to reconstruct the central emission expected for the normal galaxy population, and thus by contrast allow us to identify stars and AGN.
-4 Training and inference pipeline
-The diffusion and inpainting methods are two distinct pipelines, one for the initial training of the model and a second for the actual inpainting.
-Training the diffusion model makes use of the whole image and learns to predict what noise to remove at each timestep.
-This allows for the generation of novel examples from the trained distribution.
-4.1 Training the diffusion model
-For training, we use 6464 pixel VIS cutouts for each of the selected sources discussed in Sect.
-3.1.
-We use the training pipeline described in Dhariwal & Nichol (2021)111Available at https://github.com/openai/guided-diffusion with the loss function
-The hybrid loss is a combination of the mean squared error (MSE) and the variational lower bound (VLB).
-Using MSE ensures the output pixels of the model are numerically similar to the original image, maintaining structural and visual consistency within the generated images.
-The first term in 6 is
-where and are the true and predicted noise added for a timestep, respectively.
-The VLB loss, shown in Eqs.
-8 to 10, measures how well the diffusion process has modelled the underlying distribution of the image data, ensuring that the image as a whole is coherent and plausible for the data.
-The VLB loss has a weighting applied so as to not overpower the MSE loss:
-By utilising the Kullback-Leibler (KL) divergence (Kullback & Leibler 1951), a measure of the difference of two distributions, the loss acts as a regulariser to prevent simple memorisation and instead optimising for learning general features and patterns present in real data.
-By combining both losses, the model becomes accurate at generating realistic data at both the micro (pixel-wise) and macro (image-wise) levels.
-This is vital in ensuring the model learns the dynamics of morphology images through the diffusion process, allowing it to effectively recreate essential features through repainting.
-Throughout the rest of the paper the loss presented in Eq.
-6 is referred to as the default loss.
-Alternative loss
-Making use of standard metrics whose primary use-case is data on a consistent scale, such as 0–255 pixels, may not function optimally with the high dynamic range of VIS data.
-Specifically, since MSE is not scale invariant, having images that feature pixels that are an order of magnitude brighter, results in relative errors increasing by a square of that difference.
-This is typically a benefit as it forces outlier predictions to be severely penalised.
-However, with our data, this results in brighter images being disproportionately penalised compared to fainter ones, even if the predictions are both the same relative distance away from the true prediction.
-To address this, we applied a normalisation to by dividing each image’s error by its respective maximum pixel value, which could be from a brighter external source.
-This allows for the errors of both bright and faint images to scaled more equitably.
-The output of this loss is still combined with errors calculated in .
-Throughout the remainder of the paper, we refer to this variant of the loss as the normalised loss.
-To visualise the impact of amending the loss function, with respect to an images brightness, Fig.
-5 shows how the error from the unmodified MSE changes as the model is trained for longer.
-Each line represents the error for a collection of images in a random batch, with the increased transparency indicating the passage of time where the model had been trained on more epochs.
-The final epoch of each model is presented by a black line.
-The effect of removing point-like sources, as discussed in Sect.
-3.1, is also explored.
-An initial observation that can be made with this representation is the distribution of the number of images of a particular brightness.
-By focusing solely on the black lines, where the impact is most noticeable, each vertex on the line represents an individual image.
-The more frequent updates the line has within a range of maximum pixel values, the more images in our data set exist within that region.
-As the line has more updates when the max pixel value is less than one, there are significantly more images with fainter pixels.
-This further highlights the disproportionate penalisation of MSE as it assigns the majority of the error to a minority of the data.
-Although all four models arrive at a similar minimum loss, and begin the training process near-identically, their respective behaviour throughout training differs significantly.
-For the full data set with the default loss (top left of Fig.
-5), training prioritises brighter images, leading to some significantly low errors towards the rightmost data.
-However, this comes at the expense of fainter data, where wild fluctuations in error show that the model’s output is not very robust when it comes to precise predictions.
-This happens even when MSE might suggest low error, since pixel values below 1.0 result in minimal losses despite the predictions being relatively incorrect.
-For the normalised loss with the full data set (top right), since the penalisation is shared more fairly across pixel values, it is the density of images that has the most influence.
-Early on in training we can see that as the majority of data reduces in error, the brightest images remain at near-starting performance.
-This continues until the errors of the fainter images become so low that model is forced to prioritise the bright images.
-The key difference here is this only happens after the rest of the images are at a sufficient standard.
-The final outcome is fainter images performing well and the entire scale of images having the same maximum error.
-The two models which feature the reduced training set (bottom row) share similar dynamics during training, regardless of the loss used.
-This is largely due to many point-like sources also being some of the brightest images and therefore the scale of the data set is reduced.
-As the most problematic data are removed, the impact of disproportionate penalisation is significantly reduced, allowing the fainter sources to be the focus during training.
-After analysing the results from each of the four models, which will be discussed in Sect.
-6.2, we chose to use the model trained on the full data set using the normalised loss.
-4.2 Mask creation
-Due to the nature of the application, our aim is to mask out any suspected AGN without the need to rely on labels ahead of time.
-Each mask is positioned so that brightest pixel in the centre of the source becomes the centre of the mask.
-Although each image is centred on the source, various interactions and dynamics allow for suspected AGN to be offset from the centre.
-To allow for some deviation, a 99 pixel window around the centre of the image is used to detect the brightest pixel.
-This offset allows for the correct masking of less symmetric sources whilst limiting the likelihood of masking an adjacent source by mistake.
-The masks for each image are assigned dynamically throughout the data loading pipeline.
-Throughout this paper we use a 55 pixel mask, unless explicitly stated otherwise, centred on the calculated brightest pixel.
-A 55 pixel mask, which covers 0505, allows us to effectively cover the AGN contribution whilst minimising loss of the surrounding galaxy structure.
-Given that the VIS PSF has a full width half maximum (FWHM) of 013 (Euclid Collaboration: Mellier et al.
-2024), the AGN’s light is spread over multiple pixels.
-Taking into account the redshift of the galaxy, AGN within nearby sources () will cover 4.8 pixels, ensuring that we can fully encapsulate the core without unnecessary overlap of the host morphology.
-By retaining as much as the host galaxy information as possible for inpainting, we also reduce the risk of introducing artificial features into the reconstructed image.
-4.3 Reconstruction rescaling
-An important difference with the processing of data in this work compared to other morphology-based papers is how the data input into the network are the raw values direct from the sensor, rather than images binned to the standard 0–255 pixel scale range.
-As comparisons of individual pixels are being made on such a small mask, the relative differences become much more impactful and would likely be lost if values were binned.
-One potential issue with this setup is that the libraries used for training are optimised for images within the standard pixel range.
-Although this did not prevent the model from effective learning the distribution, it does require a rescaling of the generated output as the range is restricted to between 1 and 1 to improve training stability.
-We experimented with simply normalising the input data so that all our images were in the range 0–1; however, due to the large dynamic range of the data, the faintest pixels ended up being .
-Given the previous discussions on how MSE does not penalise smaller errors effectively, the model could not accurately recreate sources, outputting a near empty image of very low pixel values.
-Instead, to allow the comparison of the original and outputted pixel values, we make use of the fact that 99.5% of the image (4071 of 4096 pixels when using a 55 pixel mask) are shared between the input and output.
-The following scaling can be dynamically applied using each images respective values to allow the non-masked pixels to be consistent:
-where
-and
-By ensuring that the shared pixels are consistent, the inpainted pixels can also be correctly adjusted.
-5 Background noise versus diffusion noise
-It is important for clarification to address the distinctions between the definition and use of ‘noise’ in diffusion models compared to the traditional understanding of noise in astronomy-based data.
-In the context of astronomical imaging, noise refers to the random fluctuations inherent in observational data, often caused by e.g.
-background emission, electronics read-out noise, and atmospheric disturbances.
-This ‘background noise’ reduces the signal-to-noise ratio (S/N) of the image, affecting the ability to detect subtle details in any analysis.
-In contrast, the diffusion process uses an intentional integration of Gaussian noise to facilitate training of the generative model.
-The pixels of image , follow the distribution: , rather than following the distribution of the background noise.
-It is unclear whether using the background noise as the pure noise sample would cause issues for the model.
-Given the aim of the diffusion process is to denoise a noisy image so that a realistic one emerges, the original noise should be largely, if not entirely, removed.
-Given the background noise constitutes any area that is not significantly affected by a source, the model may attempt to remove it which would create an unrealistic image.
-On the other hand, it may result in less pixels needing to be denoised, optimising the models ability at learning and processing the source features.
-Such an experiment is left for future exploration.
-5.1 The effect of signal-to-noise
-In machine-learning research, the improvements of generative models are often demonstrated with data sets that are of high resolution and typically noise free.
-This provides a level of consistency across the full data set, ensuring the benchmarks reflect the model’s performance under ideal conditions.
-Whether through loss functions (as discussed in Sect.
-4) or in the practical application of the pipelines, how the model behaves on one part of the data will be similar to another.
-However, the difficulty in learning the dynamics of our data are not consistent throughout the images.
-Higher signal-to-noise ratios and lower magnitudes provide higher contrast between the source features and the background noise, enabling the model to discern these features more accurately and faster than the low S/N counterparts.
-In the context of diffusion models, where the aim throughout training is to learn the dynamics of noise at different timesteps, the varying S/N of the images could again cause inconsistencies during training.
-Figure 6 explores how this variability in data impacts the scheduled noise added to the data.
-By measuring the specific pixels that are dominantly background noise (top) and those that are significantly affect by the source (bottom), and further splitting images into S/N bins, we show the transformation from raw pixels to Gaussian noise across the entire scale of our data.
-Analysing the top row, we can see that low S/N images are more significantly affected by the introduction of noise.
-This also has an effect much earlier on in the scheduler.
-The background noise converges to the diffusion’s pure Gaussian distribution, across all S/N bins, at between 1/2 () and 3/4 () of the total noise steps.
-When focusing on the pixels dominated by the source (bottom row), high S/N images are fairly unaffected for large parts of the schedule, with the brightest images not changing significantly until half way ().
-It is not until the scheduler is within the final 250 steps that the brightest images converge to the Gaussian distribution.
-To visualise the true effect of this differing behaviour across S/N, Fig.
-7 shows a sample of images throughout different timesteps.
-Each image shown is a sample from the respective S/N bins below.
-It is clear that the noise scheduling does not perform uniformly across the images.
-To increase performance of a diffusion model, whether to recreate a data set more accurately or increase the complexity of the data you are modelling, the simplest way is to increase the number of unique denoising steps performed.
-The important part here is the requirement to be unique, as simply adding more timesteps often results in the additional timesteps (especially in the latter stages) looking too similar and therefore become just additional overhead.
-This is the exact reason Nichol & Dhariwal (2021) implemented the cosine-beta scheduler, discussed Sect.
-2.3, the original linear scheduler produced images that were pure noise too early into the process.
-The images in Fig.
-7 show that there are further optimisations that could be made to improve the scheduler for our data set.
-For optimal performance, one should expect the entire grid of images to behave like those in final 2 or 3 columns.
-A scheduler that adapts to the specific S/N of the image would be required and is left for future work.
-As our work is specifically using the reconstruction loss of inpainting as a classifier, the current scheduler is sufficient for our science goals.
+In the following sections of the paper, we will discuss the results of the diffusion model and analyse the performance of the inpainting pipeline that automatically learns to reconstruct the central emission expected for the normal galaxy population, allowing us to identify stars and AGN as outliers.
 5 Results
-To explore the ability of the model at recreating galaxy images, we apply a variety of different masks to a selection of unseen images.
-Figure 8 shows the original image on the leftmost column, followed by the regenerated images in subsequent columns.
-With each mask, shown on the top row, the white area shows the pixels that are fixed from the original image whereas the black areas are generated using the diffusion model.
-As our model’s aim is not to generate large collections of synthetic data, the variety and scientific validity of the produced galaxies are not put under high scrutiny, instead we only check that there are no obvious catastrophic mistakes.
-In the last two columns, we can see that due to the large mask and fewer pixels to guide the reconstruction, some of the images produce a visible boundary around the image where the combination of fixed pixels and generated pixels are not as seamless as the other images.
-Given the masks used for AGN detection are significantly smaller than those presented in Fig.
-8, the fraction of pixels that the inpainting is conditioned on will be much higher, providing tighter constraints on the output compared to these examples.
+Although the focus of the paper is on the model’s performance on small areas of a galaxy’s core, Repaint’s adaptability allows the pipeline to handle a variety of mask shapes and sizes.
+More general examples of the inpainting performance can be seen in Fig.
+17.
+With larger masks, fewer pixels remain for conditioning, making the output more susceptible to artifacts.
+However, Lugmayr et al.
+(2022) demonstrate that this can be mitigated by increasing the number of resamples during the Repaint loop, at the cost of longer inference times.
 5.1 Inpainting of AGN
 For our inpainting, we use a 55 pixel mask centred on the brightest pixel, following the constraints described in Sect.
-4.2.
+3.1.
 Comparing the expected pixel values with the generated outputs, Fig.
 5 shows the model’s performance.
-The assumption that the model will predict a fainter core than the original does appear to be true, with a near global bias for underpredicting as shown by the dashed line.
-The majority of sources behave in a consistent way with the exception of a smaller cluster of sources that predict more than an order of magnitude fainter than the original.
+The assumption that the model will predict a fainter core than the original does appear to be true, with a near-global bias for underpredicting, as shown by the dashed line.
+The majority of sources behave in a consistent way, with the exception of a smaller cluster of sources that predict more than an order of magnitude fainter than the original.
 This distinct cluster could provide sufficient candidates for AGN.
 The horizontal bar in the leftmost panel of Fig.
-9, where the maximum output value appears to equal one, is a side effect of Eq.
+5, where the maximum output value appears to equal one, is a side effect of Eq.
 11.
-The values themselves are not exactly one, with 1% of the data fall within of having a maximum output of one.
+The values themselves are not exactly one, with 1% of the data falling within of having a maximum output of one.
 There does exist a convergence point in Eq.
 11 , where when the maximum and minimum pixel values of an input image equal one, the output of the rescaling always equals one; however, such an image would have all its pixels valued one, and so is not present in the data.
-Although there is not a specific convergence point that explains this trend, the majority of the images that fall on this line occur when the network outputs a maximum pixel value close to one.
-However, there are many more images that have a network output close to one that are not close to this line, as well as images close to the line without a network output close to one.
-We have verified that both point-like and extended source appear within this line, therefore due to the low number of sources impacted, we leave more in-depth exploration to future work.
-To further validate the output of the model and show it is producing realistic outputs, Fig.
+Although no specific convergence point explains this trend, the majority of the images that fall on this line occur when the network outputs a maximum pixel value close to one.
+However, many more images have a network output close to one that is not close to this line, as well as images close to the line without a network output close to one.
+We have verified that both point-like and extended sources appear within this line; therefore, due to the low number of sources impacted, we leave more in-depth exploration to future work.
+To further validate the model’s output and show it is producing realistic outputs, Fig.
 19 shows how the ROC metric compares between input and output images.
-The significant overlap, especially with respect to peak of each distribution, shows that the model is accurately recreating the dynamics of the data.
+The significant overlap, especially with respect to the peak of each distribution, shows that the model is accurately recreating the dynamics of the data.
 To analyse how the error of each subset of class candidates have performed, comparisons are made between the original pixels and inpainted reconstruction for each image.
 The MSE of the pixel values are shown in Fig.
-10.
+6.
 Each subset is binned according to its magnitude, providing a distinction between the relative brightness of each image.
 There is a clear distribution difference between the respective class candidates.
-Stars perform the most consistently across magnitudes, producing the highest MSEs.
-There is an inherent trend for how high the error of the predictions are with the relative brightness of the classes, with stars and the brightest QSOs featuring the highest errors across all magnitudes.
-This is then followed by QSOs at higher magnitudes, and the lowest magnitude AGN candidates.
-Finally, the galaxy sample, along with the faintest AGN candidates produce much lower errors.
-Since these classes are candidates, it could also highlight the presence of contamination within the AGN labels.
-However, the large number of DESI-detected galaxies showing vastly reduced errors implies that the model does develop a bias for recreating galaxies with little or no AGN component.
-The second important finding is the trend between VIS magnitude and error rate across all classes.
-The brighter a source is, the higher its recorded MSE.
-This matches the initial assumptions of the method, where, as the sources increase in magnitude, differences in brightness values across galaxies become harder to differentiate.
-This leads to fainter potential AGN candidates looking more similar to a normal galaxy, causing the inpainting to reconstruct more accurately and therefore less likely to be included in our selection.
-However, the relationship between brightness and error is once again more influenced by the scale-variant nature of MSE.
-However, the lack of significant overlap for many of the selections with the galaxy class still make this metric useful.
+Brighter sources, particularly stars and luminous QSOs, produce higher errors, followed by fainter QSOs and the lowest-magnitude AGN candidates.
+Galaxies and the faintest AGN candidates show significantly lower errors, showing the model’s tendancy to reconstruct sources without strong AGN features more effectively.
+This indicates that the model has learned the anticipated bias introduced by the uneven distribution within the data set.
+While MSE is scale-dependant and influenced by the source’s VIS magnitude, the separation between the errors from the galaxy class and those of the other classes suggests it remains useful as a selection metric.
+The presence of low-error AGN candidates may also indicate contamination within the AGN labels.
+It is acknowledged that because the model is trained on the full, unlabeled galaxy population, its prior distribution does contain information about AGN, even if they are rare.
+This could, in principle, lead the model to occasionally reconstruct AGN-like features, resulting in false negatives.
+However, the strong performance in separating known AGN and QSO populations, as shown in Figure 6, suggests that the prior is sufficiently dominated by non-active galaxies for this outlier detection framework to be effective.
 To showcase how the performance of the inpainting method impacts the overall image, Fig.
-11 presents the input, output and residuals for a selection of sources, including a close up of the masked area both before and after inpainting.
-It is clear to see how bright central cores can be entirely removed with inpainting, being replaced by regions that are consistent with the surrounding pixels.
-The model is also able to recreate images that do not feature any significant core accurately.
-This shows the model has a good internal representation of galaxies and how they should behave.
+7 presents the input, output and residuals for a selection of sources, including a close-up of the masked area both before and after inpainting.
+It is clear how bright central cores can be entirely removed with inpainting and replaced by regions consistent with the surrounding pixels.
+The model can also accurately recreate images that do not feature any significant core.
+This shows that the model has a good internal representation of galaxies and how they should behave.
 5.2 Max pixel difference ratio
-Although reconstruction error allows stars and a subset of AGN to stand out from the rest, a substantial overlap between suspected AGN and DESI-classified galaxies remains, especially at higher magnitudes.
+Although reconstruction error allows stars and a subset of AGN to stand out, a substantial overlap between suspected AGN and DESI-classified galaxies remains, especially at higher magnitudes.
 Even though measured on a pixel-by-pixel scale, MSE averages away specific pixel differences, leading to less nuanced differences between input and output.
-As an example, if the reconstructed image ended up being the correct values but just shifted by one pixel, this would be penalised heavily under MSE.
-From the perspective of this paper, such a scenario would imply that the model knew a bright pixel was valid and therefore, should not be penalised.
+For example, if the reconstructed image ended up being the correct values but shifted by one pixel, this would be penalised heavily under MSE.
+From the perspective of this paper, such a scenario would represent a desirable outcome, as the model would correctly predict the presence of a bright, unresolved component, even if its exact location were shifted by a few pixels within the mask.
 Another example would be if only the brightest pixel is different in the reconstruction.
-MSE would significantly reduce the impact of this measurement, potentially scoring similar to a source that gave small errors in every pixel within the mask.
+MSE would significantly reduce the impact of this measurement, potentially scoring similarly to a source that gave minor errors in every pixel within the mask.
 To ensure such cases are not overlooked, we measure the pixel value ratio between the brightest pixel from the input mask and the brightest pixel from the reconstruction.
 As shown in Fig.
-12, the produced distributions show a much tighter spread of values for many of the AGN selections, with a skew towards higher scores.
-The majority of the galaxy subset produce a ratio less than 10, whereas a large population of suspected AGN from each selection feature a larger ratio.
+8, the produced distributions show a much tighter spread of values for many of the AGN selections, with a skew towards higher scores.
+The majority of the galaxy subset produces a ratio less than 10, whereas a large population of suspected AGN from each selection feature a larger ratio.
 Another important difference between this metric and MSE is how the results are relative, reducing the impact of magnitude on the recorded values.
-Many of the high density peaks remain consistent, even as the magnitude of the sources increase.
+Many of the high-density peaks remain consistent, even as the magnitude of the sources increases.
 This implies that this metric is more robust at detecting fainter AGN than MSE.
-The distribution also highlights the scale of the extreme values with many sources showing pixel values approaching 100 times greater than their counterparts.
+The distribution also highlights the scale of the extreme values, with many sources showing pixel values approaching 100 times greater than their counterparts.
 Exploring these large disparities can provide insights into the types of sources that are significantly sensitive to the model’s capabilities.
-Sources producing the highest values for the ratio are those where the reconstruction produces a cutout that is significantly fainter than the original image.
-As can be seen in Fig.
+Sources producing the highest values for the ratio are those where the reconstruction produces a cutout significantly fainter than the original image.
+As seen in Fig.
 20, this space is dominated by point-like objects.
-Having the sharp falloff of pixel intensity in such a small area leads to the mask covering the majority, if not the entirety of the source.
+Having the sharp falloff of pixel intensity in such a small area leads to the mask covering the majority of the source, if not the entirety.
 When the model attempts to reconstruct the missing pixels, it generates values that are coherent with the surrounding pixels.
-In these cases, the best match for empty space being more empty space.
-The behaviour of the previous examples match with the expected behaviour of the model and the initial assumptions of the paper.
-Applying inpainting to this task is expected to produce outputs that are similar to, or produce a flattened light profile of the original pixels.
-However, there exists a subset of sources that exhibit the opposite behaviour and generate a much brighter collection of pixels.
+In these cases, the point-like source contributed very little to the surrounding pixels, leading to the generation of only background noise without any source.
+The behaviour of the previous examples matches the model’s expected behaviour and the paper’s initial assumptions.
+Applying inpainting to this task is expected to produce outputs that are similar to, or produce a flattened light profile of, the original pixels.
+However, a subset of sources exists that exhibit the opposite behaviour and generate a much brighter collection of pixels.
 Figure 21 shows examples of such sources.
-When analysing the sources that exhibit this behaviour we find that they are overwhelmingly low S/N images.
-Given the difference in behaviour of these images during training, as shown in Fig.
-7, it is clear this has had an impact on the model’s behaviour with these types of images.
-Because we would like to limit this behaviour, we analyse how many of the unseen sources produce this type of output.
-The number of sources whose max pixel difference ratio is less than one for each of the four models discussed in Sect.
-4, are shown in Table 2.
-The normalised loss helps significantly in reducing the number of sources exhibiting this behaviour.
-The normalised loss trained on the full data set was selected as our applied model due to it further reducing the number of outlier predictions.
+When analysing the sources that exhibit this behaviour, we found that they have overwhelmingly low S/N images.
+Further exploration of the impact of S/N on the diffusion process is shown in Sect.
+6.1.
 Data set | Loss function | |
@@
 No point-like | 52 313 (6.17%) | 39 507 (4.66%) |
+The number of sources whose max pixel difference ratio is less than one for models trained on the two loss functions discussed in Sect.
+2.3 are shown in Table 2.
+The normalised loss significantly reduces the number of sources exhibiting this behaviour.
+The normalised loss trained on the complete data set was selected as our applied model because it further reduced the number of erroneous predictions.
 5.3 Morphology-based performance
-Given the variety of shapes and sizes of objects that will be witnessed using Euclid, it is important to verify that the model is able to handle the complexities of different galaxy morphologies.
+Given the variety of shapes and sizes of objects that will be witnessed using Euclid, it is important to verify that the model can handle the complexities of different galaxy morphologies.
 Utilising the refined morphology classifications provided from Euclid Collaboration: Walmsley et al.
@@
 22, 23, LABEL:, and 24, respectively.
-These images show how our diffusion model and the inpainting pipeline effectively adapt to different morphologies.
+These images show how our diffusion model and the inpainting pipeline can effectively adapt to different morphologies.
 For edge-on in particular, the region of bright pixels within the masked area is often narrower than the mask itself.
@@
 As seen in Fig.
-13, there is a cluster of sources in the top right.
+9, there is a cluster of sources in the top right.
 The key aspect of this cluster is its clear boundary from the DESI galaxy selection.
@@
 We create a simple decision boundary between the two clusters of points, as shown in Fig.
-13.
-Although a more fine-tuned boundary would certainly be possible, especially when incorporating a further machine-learning model, it is important to show the benefits of utilising the inpainting method through only its error metrics.
+9.
+Although a more fine-tuned boundary would be possible, especially when incorporating a secondary machine-learning model, it is important to show the benefits of utilising the inpainting method through only its error metrics.
 The equation of our linear boundary is .
@@
 This would result in a different, but possibly equally plausible, output image.
-We perform the inpainting pipeline seven separate times on the same subset of the data.
+7 samples are generated as part of a single application of the pipeline.
 To ensure that any low variance in samples is not caused by the model simply ‘remembering’ the true image from its training examples, all resampling is applied on images the model has never seen before.
-The mean and variance over a normalised by brightness MSE are shown in Fig.
-14.
+The mean and variance over a normalised-by-brightness MSE are shown in Fig.
+10.
 The top figure shows that irrespective of magnitude, the model achieves similar relative median error across all images, highlighting the impact of the normalised loss throughout training.
-However, the difference in the scheduler over images of different quality and brightness is made evident in the bottom figure.
-The consistency of generated outputs vary drastically as fainter (and therefore lower S/N) images produce widely different pixels when resampled.
-6.6 Comparing to other criteria
+However, the difference in the scheduler over images of different quality and brightness is evident in the bottom figure.
+The consistency of generated outputs varies drastically as fainter (and therefore typically lower S/N) images produce widely different pixels when resampled.
+5.6 Comparing to other classifiers
 It is important to compare our AGN and QSO predictions to more traditional methods that are also able to utilise only the image data.
-Given the assumption of our model that sources containing an AGN will feature a bright central component with a sharp falloff in the galaxy’s light profile, comparing to Srsic flux and point spread function (PSF) flux measurements is crucial because they capture the key concepts of our assumptions.
-The Srsic flux helps quantify galaxy light distribution and differentiate an AGN’s central brightness from the galaxy’s structural brightness.
-PSF flux isolates the galaxy’s core, capturing point-like behaviour.
-The distribution of sources with these parameters is shown in Fig.
-15.
-The equation of this linear boundary is .
+Srsic and PSF Flux Measurements
+To benchmark our non-parametric inpainting method against a more traditional approach, we construct a classifier based on parametric modelling of galaxy light profiles.
+We utilise morphological parameters sourced directly from the Euclid MER Q1 catalogue (Euclid Collaboration: Romelli et al.
+2025).
+Specifically, we use the provided flux_vis_sersic_euclid and flux_vis_psf_euclid measurements.
+These parameters are derived by the MER pipeline using the SourceExtractor++ software (Bertin et al.
+2020; Kümmel et al.
+2020).
+For each detected source, the pipeline performs two separate and independent model fits.
+The flux_vis_psf_euclid value represents the total flux obtained by fitting a model of the instrumental PSF to the source image.
+The flux_vis_sersic_euclid value represents the total flux from fitting a two-dimensional Srsic profile, which is designed to model the extended light distribution of a galaxy.
+The underlying assumption for a classifier based on these parameters is that sources with a dominant, unresolved central component (such as a QSO or a Type-1 AGN) will be well-described by a PSF model, resulting in a high flux_vis_psf_euclid value.
+Conversely, normal resolved galaxies will be better described by the extended Srsic profile, leading to a high flux_vis_sersic_euclid value.
+By comparing the relative strength of these two flux measurements (as shown in Fig.
+15), we can construct a simple linear classifier to distinguish between point-like and extended sources.
+This baseline method is computationally inexpensive as it relies on pre-computed catalogue values, but it is fundamentally limited to this simple morphological distinction and may struggle with identifying AGN in complex host galaxies or at lower relative luminosities.
+The distribution of sources with these parameters is shown in Fig. 11. The equation of this linear boundary is .
+ROC Metric
 Our second comparison is with the original rate of change metric used in Sect.
-3.3.
-As we were able to see a clear distinction between the galaxy class and the others, it is important to verify how such a metric performs as a classifier.
-This is especially true given such a metric can be generated from the raw image, with no machine-learning model required.
+4.3.
+As a clear distinction between the galaxy class and the others is visible, it is important to verify how such a metric performs as a classifier.
+This is especially true given that such a metric can be generated from the raw image, and no machine-learning model or precomputed values are required.
 The boundary for this classifier is .
-The test set scores for each respective selection is shown in Fig.
-16, with the collated results for stars, galaxies, AGN and QSOs shown in Fig.
+Performance against other candidates
+Given the variety and differences in the selection methods, there is an inherent incompleteness and potential for contamination.
+Therefore our analysis focuses primarily on achieving sufficiently high recall.
+For an initial candidate selection method, such as the one proposed here, the primary goal is often to compile a large, inclusive sample for potential follow-up observations or more detailed multi-wavelength analysis.
+In this context, high recall is typically prioritised over high precision, which measures the purity of the selected sample.
+The scientific cost of missing a potentially interesting source (a false negative) is often considered higher than the operational cost of including a contaminant (a false positive) that can be filtered out in subsequent analysis stages.
+Therefore, our primary goal is to demonstrate the capability of our model to achieve high recall across a wide variety of AGN and QSO candidates which have been selected via different, more conventional methods and physical mechanisms.
+Precision and Recall are defined as follows
+For a given classification task (e.g., identifying DESI QSO objects):
+TP (true positives) is the number of objects that belong to the class and are correctly predicted as belonging to it by the model.
+FP (false positives) is the number of objects that do not belong to the class but are incorrectly predicted as belonging to it by the model.
+FN (false negatives) is the number of objects that belong to the class but are not predicted as belonging to it by the model.
+The test set scores for each respective selection are shown in Fig.
+12, with the collated results for stars, galaxies, AGN and QSOs shown in Fig.
 25.
-All predictions are on images that were not used during training, nor used in the creation of any of the decision boundaries for the classifiers.
-Given the variety and differences in the selection methods, there is an inherent incompleteness and potential for contamination, therefore our analysis focuses primarily on achieving sufficiently high recall.
-High recall ensures that our method is capable of identifying the broad range of AGN and QSO sources that are detectable through existing techniques.
-A high recall metric indicates that our classifier effectively captures the variety of sources each respective selection criterion is designed to detect.
-While precision is also a consideration, our primary metric of success in this context is the diffusion model’s recall.
-Having the ability to recover a large fraction of known AGN and QSO types, across various selection methods, validates the robustness and comprehensiveness of our classifier.
-Due to differing constraints for each of the selections, we split the analysis to allow for a more representative comparison.
+All predictions are made from images that were not used during training nor used in the creation of any of the decision boundaries for the classifiers.
+Due to differing constraints for each selection, we split the analysis to allow for a more representative comparison.
 Galaxies
@@
 2024) along with the contaminants removed from the original SPECTYPE=GALAXY selections.
-The predictions from each classifier for whether a source is a normal galaxy is simply the inverse of the boundary’s original predictions.
+The predictions from each classifier for whether a source is a normal galaxy are simply the inverse of the boundary’s original predictions.
 As Fig.
 12 shows, all classifiers capture nearly all galaxies across every magnitude bin.
+This behaviour is expected, as the selection criteria used by each classifier are explicitly designed to favour AGN-like features, thereby excluding typical galaxies by construction.
 The precision scores also match extremely closely between classifiers, indicating similar contaminants proving difficult, especially for the brighter sources.
 Such contaminants can be seen in Figs.
-13 and 15, where stars and DESI-selected QSO fall outside the boundaries.
-The ability to separate normal galaxies from other extragalactic or stellar objects has benefits for both groups of data.
-Being able to create a pure galaxy sample is vital for the precise measurements needed for cosmology, as well as to further our understand of dark matter and dark energy, a primary object of the Euclid mission (Euclid Collaboration: Mellier et al.
-2024).
+9 and 11, where stars and DESI-selected QSO fall outside the boundaries.
+The ability to separate normal galaxies from other extragalactic or stellar objects benefits both groups of data.
+Being able to create a pure galaxy sample is vital for the precise measurements needed for cosmology, as well as to further our understanding of dark matter and dark energy, a primary object of the Euclid mission (Euclid Collaboration: Mellier et al.
+2025).
 Stars
@@
 The near-perfect recall shows that our classifier is able to capture stars well.
-This performance is beneficial when attempting to capture a high purity galaxy sample as discussed in the previous section.
+This performance is beneficial when attempting to capture a high-purity galaxy sample, as discussed in the previous section.
 However, as other works such as Euclid Collaboration: Matamoro Zatarain et al.
 (2025) have shown, there are metrics to mitigate the contamination of stars.
-This is shown in practise by the high precision, highlighting how we are able to isolate this star selection from other point-like sources.
-However, a significant number of stars are likely to remain in our sample.
-Therefore, for all subsequent selections, we apply the cut phz_star_prob_phz_class, to predictions of all classifiers to minimise the contamination of stellar sources.
+As a significant number of stars are likely to remain in our sample, for all subsequent selections, we apply the cut phz_star_prob_phz_class, to predictions of all classifiers to minimise the contamination of stellar sources.
 Euclid-detected QSOs
@@
 2024).
-This subset consider only sources with MUMAX_MINUS_MAG , a parameter that indicates point-likeness (Euclid Collaboration: Romelli et al.
+This subset considers only sources with MUMAX_MINUS_MAG , a parameter that indicates point-likeness (Euclid Collaboration: Romelli et al.
 2025).
+As these selections have yet to go through a follow-up validation, they remain candidates and as such should not be seen as complete ground truth.
 For the _ and _ selections, we capture around 90% of the candidate QSOs.
@@
 As these sources are assumed point-like, applying the above MUMAX_MINUS_MAG limit would be reasonable.
-However, due to the improvements in spatial resolution from DESI, at about , to Euclid’s now 02, some sources that appeared point-like through DESI, now appear extended.
+However, due to the improvements in spatial resolution from DESI, at about , to Euclid’s now 02, some sources that appeared point-like through DESI now appear extended.
 Therefore, to ensure all potential sources are found, we do not apply the restriction to this subset.
-This selection provides the first evidence of significant differences within the candidates for each classifier.
-The inpainting classifier is able to capture a significant amount of the selection compared to the other classifiers.
+This selection provides the first evidence of significant differences in the suggested candidates for each classifier.
+The inpainting classifier is able to capture a significantly higher selection of lower magnitude sources compared to the other classifiers.
 Throughout all magnitude bands, the PSF and Srsic flux classifier has worse recall.
-Whereas the diffusion model has higher coverage over the brightest sources (lowest magnitudes), the ROC classifer matched more candidates in the fainter sources.
+The ROC classifier matched more candidates in the fainter sources.
 Gaia-detected QSOs
@@
 The subset consists of positive instances (qso_quaia, MZ25), with negative instances from Gaia-detected galaxies (in_galaxy_candidates_gaia, MZ25) and stars (star_candidate_gaia, MZ25).
-We once again apply the MUMAX_MINUS_MAG restriction to the sources.
-After the initial magnitude bin, our classifier was able to match with a high proportion of the Quaia-selected QSOs.
-The other two classifiers achieved equal recall scores, but in the faintest detected sources our classifier was able to match with an additional 20% of sources.
+We again apply the MUMAX_MINUS_MAG restriction to the sources.
+After the initial magnitude bin, our classifier matched with a high proportion of the Quaia-selected QSOs.
+The other two classifiers achieved equal recall scores, but in the faintest detected sources, our classifier was able to match with an additional 20% of sources.
 The particularly poor precision for this selection across all the classifiers is due to the considerable imbalance between positive and negative predictions.
-With only 0.5% of the Gaia-detected sources being Quaia candidates, selecting more than just the candidates is likely, significantly impacting the precision.
+With only 0.5% of the Gaia-detected sources being Quaia candidates, selecting more than just the candidates is very likely, significantly impacting the precision.
 AllWISE R90 AGN
@@
 Using the selection criteria derived by Assef et al.
-(2018), we compare our candidates to the 90% reliability (R90) selection which is a high purity AGN sample.
+(2018), we compare our candidates to the 90% reliability (R90) selection, a high purity AGN sample.
 The Assef et al.
@@
 Stars | 0.99 | 0.99 | 1.00 | 0.96 | 0.97 | 0.96 | 0.50 | 0.51 | 0.50 | ||||
-| 0.97 | 0.98 | 1.00 | 0.84 | 0.85 | 0.84 | 0.51 | 0.53 | 0.50 | |||||
+0.97 | 0.98 | 1.00 | 0.84 | 0.85 | 0.84 | 0.51 | 0.53 | 0.50 | |||||
 Galaxy DESI | 1.00 | 1.00 | 1.00 | 0.74 | 0.73 | 0.73 | 0.64 | 0.63 | 0.64 | ||||
-| 0.98 | 1.00 | 0.99 | 0.90 | 0.90 | 0.91 | 0.73 | 0.72 | 0.75 | |||||
+0.98 | 1.00 | 0.99 | 0.90 | 0.90 | 0.91 | 0.73 | 0.72 | 0.75 | |||||
 QSO DESI | 0.77 | 0.57 | 0.63 | 0.18 | 0.14 | 0.15 | 0.84 | 0.75 | 0.77 | ||||
-| 0.90 | 0.84 | 0.96 | 0.59 | 0.67 | 0.62 | 0.93 | 0.91 | 0.96 | |||||
+0.90 | 0.84 | 0.96 | 0.59 | 0.67 | 0.62 | 0.93 | 0.91 | 0.96 | |||||
 QSO QUAIA | 1.00 | 0.98 | 0.98 | 0.01 | 0.01 | 0.01 | 0.59 | 0.58 | 0.58 | ||||
-| 1.00 | 0.80 | 0.80 | 0.01 | 0.01 | 0.01 | 0.51 | 0.41 | 0.40 | |||||
-| AGN Xray | 0.64 | 0.48 | 0.61 | 0.74 | 0.73 | 0.74 | 0.63 | 0.59 | 0.62 | ||||
-| 0.51 | 0.38 | 0.53 | 0.86 | 0.86 | 0.87 | 0.53 | 0.51 | 0.55 | |||||
+1.00 | 0.80 | 0.80 | 0.01 | 0.01 | 0.01 | 0.51 | 0.41 | 0.40 | |||||
+AGN X-ray | 0.64 | 0.48 | 0.61 | 0.74 | 0.73 | 0.74 | 0.63 | 0.59 | 0.62 | ||||
+0.51 | 0.38 | 0.53 | 0.86 | 0.86 | 0.87 | 0.53 | 0.51 | 0.55 | |||||
 AGN R90 AllWISE | 0.68 | 0.57 | 0.63 | 0.02 | 0.02 | 0.02 | 0.68 | 0.63 | 0.66 | ||||
-| 0.59 | 0.45 | 0.60 | 0.03 | 0.02 | 0.03 | 0.75 | 0.68 | 0.75 | |||||
+0.59 | 0.45 | 0.60 | 0.03 | 0.02 | 0.03 | 0.75 | 0.68 | 0.75 | |||||
 QSO _ | 1.00 | 1.00 | 1.00 | 0.02 | 0.02 | 0.02 | 0.50 | 0.50 | 0.50 | ||||
-| 0.94 | 0.92 | 1.00 | 0.05 | 0.05 | 0.05 | 0.49 | 0.47 | 0.50 | |||||
+0.94 | 0.92 | 1.00 | 0.05 | 0.05 | 0.05 | 0.49 | 0.47 | 0.50 | |||||
 QSO _ | 1.00 | 1.00 | 1.00 | 0.01 | 0.01 | 0.01 | 0.50 | 0.50 | 0.50 | ||||
-| 0.95 | 0.91 | 1.00 | 0.04 | 0.04 | 0.04 | 0.50 | 0.46 | 0.50 | |||||
+0.95 | 0.91 | 1.00 | 0.04 | 0.04 | 0.04 | 0.50 | 0.46 | 0.50 | |||||
 QSO Bisigello24-A | 1.00 | 0.99 | 1.00 | 0.02 | 0.02 | 0.02 | 0.50 | 0.50 | 0.50 | ||||
-| 0.96 | 0.97 | 1.00 | 0.03 | 0.03 | 0.03 | 0.50 | 0.50 | 0.50 | |||||
-Xray-selected AGN
+0.96 | 0.97 | 1.00 | 0.03 | 0.03 | 0.03 | 0.50 | 0.50 | 0.50 |
+X-ray-selected AGN
 We compare our candidates to the X-ray selection from the work by Euclid Collaboration: Roster et al.
@@
 Only sources with detection in X-rays are included in the results.
-The main difference with the results of the X-ray-detected AGN is how the precision outperforms the recall across nearly every magnitude bin.
-This is due to low number of X-ray sources compared to the other selections.
+The main difference with the results of the X-ray-detected AGN and the other selections is that the precision outperforms the recall across nearly every magnitude bin.
+This is due to the low number of X-ray sources available.
 The recall performance between the diffusion model and the ROC classifier is very similar, with the PSF and Srsic flux classifier again struggling to match as many candidates as the other two.
-Each of the selection criteria we compare with captures distinct subsets of the AGN and QSO populations, influenced by specific observational biases and methodological constraints.
-Therefore, it is only when collating these selections and by identifying trends and overlaps in the labels, that we can be confident that we sufficiently cover the spectrum of possible AGN.
+Each of the selection criteria we compare captures distinct subsets of the AGN and QSO populations, influenced by specific observational biases and methodological constraints.
+Therefore, it is only when collating these selections and identifying trends and overlaps in the labels that we can be confident that we sufficiently cover the spectrum of possible AGN.
 Table 3 summarises the scores shown in Fig.
-16.
-We combine the results into two magnitude bins 1820 and 2022.5, separating into brighter and fainter sources respectively.
+12.
+We combine the results into two magnitude bins 18–20 and 20–22.5, separating into brighter and fainter sources, respectively.
 Due to similar scores in many of the selections, we highlight in bold any scores that outperform the other classifiers by at least 0.05.
@@
 Given our model’s ability to classify suspected AGN, it is not unreasonable for the reader to assume that the residuals of the repainting process and the output image can independently represent the AGN and galaxy components.
-Although utilising inpainting for decomposition is a worthwhile future research direction, the validity of the residuals to accurately represent the AGN component for further scientific analysis have not been tested and as such cannot be used.
-More significant exploration of which mask to use will also be vital for such a use-case.
+Although utilising inpainting for decomposition is a worthwhile future research direction, the validity of the residuals to accurately represent the AGN component for further scientific analysis has not been tested and, as such, cannot be used.
+A more significant exploration of which mask to use will also be vital for such a use case.
 A mask that adapts its size and shape for each input image will likely be necessary.
@@
 (2025), hereafter called MB25, whose method provides AGN fraction () measurements that we use to analyse our AGN candidates.
-The subset of Q1 used within this work and the subset tested in MB25, feature an overlap of 425 781 sources.
+The subset of the sources from the Q1 catalogue (Euclid Collaboration: Romelli et al.
+2025) used within this work and the subset tested in MB25 feature an overlap of 425 781 sources.
 The confusion matrix for AGN predictions between the diffusion classifier and sources with an (AGN candidate for MB25) are shown in Table 4.
 The majority of sources () are in agreement by both methods not to be AGN candidates.
-With the remaining sources, only are in agreement, with over five times more AGN candidates selected by MB25.
+With the remaining sources, only agree, with over five times more AGN candidates selected by MB25.
 When comparing all the AGN candidates selected using their respective subsets of data, this number reduces down to 3.5 times more candidates selected by MB25 than the diffusion model (Table 5).
 The left panel of Fig.
-17, shows the AGN candidates selected by the diffusion, but predicted to have low AGN fraction by the model in MB25.
-Whereas the right panel shows the AGN fraction of agreed sources between the two methods.
+13 shows the AGN candidates selected by the diffusion classifier but predicted to have low AGN fraction by the model in MB25.
+The right panel shows the AGN fraction of sources that are agreed upon between the two methods.
 Fig.
@@
 Comparing the predictions for each of the Euclid-based selections in Fig.
-18, we can see that our diffusion model covers areas of the colour space that the photometry-based selections do not.
+14, we can see that our diffusion model covers areas of the colour space that the photometry-based selections do not.
 Being able to select both bluer galaxies, as captured by the photometry selections (MZ25; Euclid Collaboration: Bisigello et al.
 2024), and the much redder sources captured by MB25, shows the versatility of the method, highlighting its effectiveness across different galaxy populations.
-7 Summary and conclusions
-In this paper, we have used a diffusion-based inpainting model to identify AGN and QSOs using VIS images from the Q1 data (Euclid Quick Release Q1 2025) from the Euclid telescope.
-Our approach focuses on inpainting the brightest pixels within the centre of a galaxy, and employs a novel thresholding approach based on the reconstruction errors to differentiate AGN and QSOs from the rest of the galaxy population.
-Using only VIS images, our method can generate a large, reliable sample of optically-selected AGN, without the requirement of AGN labels ahead of time.
-Utilising a standard training and inference pipeline with minimal modifications for astronomical data, our method demonstrates high recall and significant overlap with AGN selections obtained through more traditional methods, across various wavelengths including optical, near-infrared (NIR), mid-infrared, and Xray.
+6 Discussion
+6.1 The effect of signal-to-noise
+In machine-learning research, the improvements of generative models are often demonstrated with data sets that are of high resolution and typically noise-free.
+This provides a level of consistency across the full data set, ensuring the benchmarks reflect the model’s performance under ideal conditions.
+Whether through loss functions (as discussed in Sect.
+2.3) or in the practical application of the pipelines, the model’s response to one part of the data will be similar to another.
+However, the difficulty in learning the dynamics of our data is not consistent throughout the images.
+Higher signal-to-noise ratios and lower magnitudes provide higher contrast between the source features and the background noise, enabling the model to discern these features more accurately and faster than the low S/N counterparts.
+In the context of diffusion models, where the aim throughout training is to learn the dynamics of the introduced noise at different timesteps, the varying S/N of the images could again cause inconsistencies.
+This can be seen in both the classifier results shown in Fig.
+12, where the fainter objects typically have lower recall, as well as the increased variance during inpainting of fainter images shown in Fig.
+10.
+To explore how the noise schedule impacts the Euclid images, Figure 15 shows the variability of pixel values as they progress through timesteps.
+We take a row and column of pixels (a slice) through each image, bin them by S/N, and compute the median pixel value at each index within the slice.
+This allows us to observe how the values change across the image.
+Since pixels affected by the source have much higher values than those near the edges of the image, Figure 15 compares slices through the outer regions (top) with slices through the centre of the image (bottom).
+In the top row, we see that low S/N images are more significantly affected by the introduction of noise, and this effect appears earlier in the scheduler.
+The background noise converges to the Gaussian distribution across all S/N bins between halfway () and three-quarters () of the total noise steps.
+In contrast, for pixels dominated by the source (bottom row), high S/N images remain relatively unaffected through much of the schedule.
+The brightest images do not show noticeable changes until around , and only converge to the Gaussian distribution within the final 250 steps.
+It is evident that both the image’s S/N level and the position of the pixels within the image determine the number of timesteps needed for the noise to resemble the target Gaussian distribution.
+To visualise the true effect of this differing behaviour, Fig.
+16 shows a sample of images throughout different timesteps.
+Each image shown is a sample from the respective S/N bins below.
+It is clear that the noise scheduling does not perform uniformly across the images.
+To improve the performance of a diffusion model, whether to better recreate a data set or to model more complex data, the simplest approach is to increase the number of unique denoising steps.
+The emphasis here is on uniqueness, as simply adding more timesteps can lead to redundancy, especially in the later stages where successive steps become too similar, add little value and increase overhead.
+This is precisely why Nichol & Dhariwal (2021) introduced the cosine-beta scheduler (discussed in Sect.
+2.4); the original linear scheduler caused images to degrade into pure noise too early into the process.
+The images in Fig.
+16 show that for a large fraction of our data set, the model spends the majority of its time and computational effort learning the trivial task of denoising pure Gaussian noise, which provides no useful learning signal for galaxy morphology.
+This is a fundamental inefficiency in the implementation of diffusion-based models that not only affects Euclid images, but will apply to any scientific data set that exhibit varying levels of S/N.
+Although this has implications in regards to the efficiency and optimisation of the pipeline, it highlights the bigger issue of the data quality and variety assumptions implied when using off-the-shelf deep learning methods to the complex, heterogeneous data typical in large astronomical surveys and the wider scientific community.
+While a detailed implementation is beyond the scope of this work, this finding highlights a critical direction for future research: the development of adaptive noise schedulers that are conditioned on image properties like S/N.
+Such an advance would significantly improve both the efficiency and performance of diffusion models in scientific applications.
 6.2 Considerations on training and inference costs
@@
 This memory constraint is due to the need to store and iteratively manipulate large intermediate tensors during inference.
-Whilst training, although more computationally expensive, has a more structured process where gradients are computed and accumulated in a predictable manner.
-With an expanded compute resource (4 Nvidia H100 GPUs), we are able to inpaint just over 2.5 million images per day.
+Training, although more computationally expensive, has a more structured process where gradients are computed and accumulated in a predictable manner.
+With an expanded compute resource (4 Nvidia H100 GPUs), we can inpaint just over 2.5 million images per day.
 As we have shown in Sects.
-4 and 2.3, our model architecture and hyperparameters remain largely unmodified from their standard configurations.
-Optimisations tailored explicitly for astronomical-based data, especially in regards to the noise scheduler, will significantly reduce the number of iterations and inference time required to process each image.
+3 and 2.4, our model architecture and hyperparameters remain largely unmodified from their standard configurations.
+Optimisations tailored explicitly for astronomical-based data, especially in regards to the noise scheduler discussed in Sect.
+6.1, will significantly reduce the number of iterations and inference time required to process each image.
 For example, Fig.
@@
 5.6 show that although our predictions have a significant overlap with many of the selections presented, it also highlights that the ROC classifier produces similar overlap performance.
-Given this boundary is formed by a simple calculation on the data, rather than requiring a complex model, it may present itself as the preferred method.
-However, it must be reiterated that the current pipeline has had minimum optimisations for the data and the classification task.
-Therefore, the ROC classifier serves as a fundamental baseline for comparison in our study.
-Future enhancements to the model are expected to not only improve the accuracy of identifying AGN candidates but also refine our reconstruction error maps for better analysis of prospective AGN components.
+Given that this boundary is formed by a simple calculation from the data, rather than requiring a complex model and therefore near-zero overhead to calculate, it may present itself as the preferred method.
+However, the ROC metric is static and its performance is fixed by its mathematical definition.
+It must be reiterated that the current pipeline for the diffusion model has had minimum optimisations for the data and the classification task and so can be expected to improve with larger training sets, more advanced network architectures, or the future incorporation of multi-wavelength data as conditioning information.
+The ROC classifier is also limited to outputting a single number.
+The generative capabilities of the diffusion model, where it outputs a reconstructed image, provides the potential for a much richer set of scientific applications beyond simple classification such as decomposition as was discussed in Sect.
+5.7.
+Therefore, although it has shown good performance, its limited approach restricts the ROC classifier to only serve as a fundamental baseline for comparison in our study.
+In regards to the PSF and Srsic flux classifier, although there are no computational costs to the user, traditional parametric fitting is a serial, per-object process that becomes a bottleneck for survey-scale datasets (Tuccillo et al.
+2018).
+Having parallelisable and scalable alternative methods, even if they feature a large one-time training cost, allow for a strategic investment as we cover more of the extragalactic sky.
+For AGN identification specifically, the simple parametric models may not be well suited to handle complex or merged morphologies.
+This could explain the worse AGN-based recall seen for this classifier in Fig.
+12.
+Future enhancements to the diffusion model are expected to not only improve the accuracy of identifying AGN candidates but also refine our reconstruction error maps for better analysis of prospective AGN components.
 The other classifiers do not offer the same potential for improvement, making the diffusion method the most promising moving forward.
-Our key contributions and conclusions include the following
-Compared to traditional colour-selection criteria, the approach is attaining high coverage in AGN candidates while achieving near-perfect separation of galaxies not featuring an AGN component.
-Reliably separating these different sources allows for creation of subsets with high purity which are vital for the cosmological objectives of the Euclid mission.
+7 Conclusion
+In this paper, we have used a diffusion-based inpainting model to identify AGN and QSOs using VIS images from the Q1 data (Euclid Quick Release Q1 2025) from the Euclid telescope.
+Our approach focuses on inpainting the brightest pixels within the centre of a galaxy and employs a novel thresholding approach based on the reconstruction errors to differentiate AGN and QSOs from the rest of the galaxy population.
+Using only VIS images, our method can generate a large, reliable sample of optically-selected AGN without requiring AGN labels ahead of time.
+Utilising a standard training and inference pipeline with minimal modifications for astronomical data, our method demonstrates high recall and significant overlap with AGN candidates obtained through more traditional methods, across various wavelengths including optical, near-infrared (NIR), mid-infrared, and X-ray.
+Our key contributions and conclusions include the following:
+Compared to traditional colour-selection criteria, the approach attains high coverage in AGN candidates while achieving near-perfect separation of galaxies not featuring an AGN component.
+Reliably separating these different sources allows for the creation of subsets with high purity, which are vital for the cosmological objectives of the Euclid mission.
 By demonstrating that VIS images alone can generate reliable and extensive samples of AGN and QSOs, our method showcases the impressive capabilities of Euclid’s imaging.
-Given the future of the mission will cover , the ability for AGN identification using only VIS allows for the possibility of some science aims to be met without the immediate need for integration of multiple data sources.
-This can significantly reduce time and complexity for processing and analysing the data.
+Given that the future of the mission will cover , the ability for AGN identification using only VIS allows for the possibility of some science aims to be met without the immediate need for integrating multiple data sources.
+This can significantly reduce the time and complexity of processing and analysing the data.
 The application of diffusion models with minimal adaptations for astronomical imaging highlights the potential for the use of more advanced machine-learning tools and pipelines for future Euclid data, as well as other scientific data sets.
@@
 Although our model performs well without significant optimisation for our use case, we provide insights into the behaviour of the model given the challenges presented with astronomical imaging.
-We show how the assumptions of machine-learning literature can result in non-optimal training, especially when it comes to the noise scheduler and training losses.
-By highlighting these inefficiencies, we can see where progress can be made in both astronomy and machine-learning domains.
+We show how the assumptions of machine-learning literature, particularly regarding fixed noise schedules, can result in non-optimal training on data with a high dynamic range and large variations in signal-to-noise.
+By highlighting these inefficiencies, we identify a clear direction for future progress in both astronomy and machine-learning domains through the development of adaptive, data-aware training strategies.
 Because this is the first use of these pipelines for such an application, we have routinely provided guidance on improvements and next steps to refine this work.
-Enhancing the model’s precision, improving the efficiency of both training and inference, as well as expanding its capabilities, will only increase its impact.
+Enhancing the model’s precision, improving the efficiency of both training and inference and expanding its capabilities will only increase its impact.
 In conclusion, our results demonstrate the feasibility and effectiveness of using diffusion-based inpainting for AGN and QSO identification.
-The ability to use VIS images exclusively for selecting these candidates, provides a potential strategic flexibility to the Euclid mission whilst aiding in reducing the reliance on complex multispectral data integration.
+The ability to use VIS images exclusively for selecting these candidates provides a potential strategic flexibility to the Euclid mission while reducing reliance on complex multispectral data integration.
 By proving the utility of applying our pipelines to the task of AGN identification, we believe it can be expanded and can significantly contribute to a broader range of astronomical applications, such as identifying gravitational lensing or finding transient objects.
@@
 […]
-- Adam et al. (2022) Adam, A., Coogan, A., Malkin, N., et al. 2022, arXiv preprint arXiv:2211.03812
-- Adam et al. (2023) Adam, A., Stone, C., Bottrell, C., et al. 2023, arXiv preprint arXiv:2311.18002
-- Assef et al. (2018) Assef, R. J., Stern, D., Noirot, G., et al. 2018, ApJS, 234, 23
-- Davis et al. (2018) Davis, B. L., Graham, A. W., & Cameron, E. 2018, ApJ, 869, 113
-- Davis et al. (2019) Davis, B. L., Graham, A. W., & Cameron, E. 2019, ApJ, 873, 85
-- DESI Collaboration et al. (2024) DESI Collaboration, Adame, A. G., Aguilar, J., et al. 2024, AJ, 168, 58
-- Dhariwal & Nichol (2021) Dhariwal, P. & Nichol, A. 2021, Advances in neural information processing systems, 34, 8780
-- Dressler (1989) Dressler, A. 1989, in IAU Symposium, Vol. 134, Active Galactic Nuclei, ed. D. E. Osterbrock & J. S. Miller, 217
-- Euclid Collaboration: Bisigello et al. (2024) Euclid Collaboration: Bisigello, L., Massimo, M., Tortora, C., et al. 2024, A&A, 691, A1
-- Euclid Collaboration: Cropper et al. (2024) Euclid Collaboration: Cropper, M., Al Bahlawan, A., Amiaux, J., et al. 2024, A&A, accepted, arXiv:2405.13492
-- Euclid Collaboration: Margalef-Bentabol et al. (2025) Euclid Collaboration: Margalef-Bentabol, B., Wang, L., La Marca, A., et al. 2025, A&A, submitted
-- Euclid Collaboration: Matamoro Zatarain et al. (2025) Euclid Collaboration: Matamoro Zatarain, T., Fotopoulou, S., Ricci, F., et al. 2025, A&A, submitted
-- Euclid Collaboration: Mellier et al. (2024) Euclid Collaboration: Mellier, Y., Abdurro’uf, Acevedo Barroso, J., et al. 2024, A&A, accepted, arXiv:2405.13491
-- Euclid Collaboration: Romelli et al. (2025) Euclid Collaboration: Romelli, E., Kümmel, M., Dole, H., et al. 2025, A&A, submitted
-- Euclid Collaboration: Roster et al. (2025) Euclid Collaboration: Roster, W., Salvato, M., Buchner, J., et al. 2025, A&A, submitted
-- Euclid Collaboration: Walmsley et al. (2025) Euclid Collaboration: Walmsley, M., Huertas-Company, M., Quilley, L., et al. 2025, A&A, submitted
-- Euclid Quick Release Q1 (2025) Euclid Quick Release Q1. 2025, https://doi.org/10.57780/esa-2853f3b
-- Ferrarese & Merritt (2000) Ferrarese, L. & Merritt, D. 2000, ApJ, 539, L9
-- Gaia Collaboration: Bailer-Jones et al. (2023) Gaia Collaboration: Bailer-Jones, C. A. L., Teyssier, D., Delchambre, L., et al. 2023, A&A, 674, A41
-- Gaia Collaboration: Vallenari et al. (2023) Gaia Collaboration: Vallenari, A., Brown, A. G. A., Prusti, T., et al. 2023, A&A, 674, A1
-- Goodfellow et al. (2014) Goodfellow, I., Pouget-Abadie, J., Mirza, M., et al. 2014, Advances in neural information processing systems, 27
-- Goodfellow et al. (2020) Goodfellow, I., Pouget-Abadie, J., Mirza, M., et al. 2020, Communications of the ACM, 63, 139
-- Harrison & Ramos Almeida (2024) Harrison, C. M. & Ramos Almeida, C. 2024, Galaxies, 12, 17
-- Ho et al. (2020) Ho, J., Jain, A., & Abbeel, P. 2020, Advances in neural information processing systems, 33, 6840
-- Karras et al. (2020) Karras, T., Laine, S., Aittala, M., et al. 2020, in IEEE/CVF conference on computer vision and pattern recognition (IEEE), 8110–8119
-- Kingma (2013) Kingma, D. P. 2013, arXiv preprint arXiv:1312.6114
-- Kormendy & Richstone (1995) Kormendy, J. & Richstone, D. 1995, ARA&A, 33, 581
-- Kullback & Leibler (1951) Kullback, S. & Leibler, R. A. 1951, The annals of mathematical statistics, 22, 79
-- Lugmayr et al. (2022) Lugmayr, A., Danelljan, M., Romero, A., et al. 2022, in IEEE/CVF conference on computer vision and pattern recognition (IEEE), 11461–11471
-- Lupton et al. (2004) Lupton, R., Blanton, M. R., Fekete, G., et al. 2004, PASP, 116, 133
-- Magorrian et al. (1998) Magorrian, J., Tremaine, S., Richstone, D., et al. 1998, AJ, 115, 2285
-- Nichol & Dhariwal (2021) Nichol, A.
-Q.
-& Dhariwal, P.
-2021, in Proceedings of Machine Learning Research, Vol.
-139, Proceedings of the 38th International Conference on Machine Learning, ed.
-M.
-Meila & T.
-Zhang (PMLR), 8162–8171
-- Pathak et al. (2016) Pathak, D., Krahenbuhl, P., Donahue, J., Darrell, T., & Efros, A. A. 2016, in IEEE conference on computer vision and pattern recognition (IEEE), 2536–2544
-- Radford et al.
-(2021) Radford, A., Kim, J.
-W., Hallacy, C., et al.
-2021, in Proceedings of Machine Learning Research, Vol.
-139, Proceedings of the 38th International Conference on Machine Learning, ed.
-M.
-Meila & T.
-Zhang (PMLR), 8748–8763
-- Ramesh et al. (2022) Ramesh, A., Dhariwal, P., Nichol, A., Chu, C., & Chen, M. 2022, arXiv e-prints, arXiv:2204.06125
-- Remy et al. (2023) Remy, B., Lanusse, F., Jeffrey, N., et al. 2023, A&A, 672, A51
-- Saharia et al.
-(2022) Saharia, C., Chan, W., Saxena, S., et al.
-2022, in Advances in Neural Information Processing Systems, ed.
-S.
-Koyejo, S.
-Mohamed, A.
-Agarwal, D.
-Belgrave, K.
-Cho, & A.
-Oh, Vol.
-35 (Curran Associates, Inc.), 36479–36494
-- Sahu et al. (2019) Sahu, N., Graham, A. W., & Davis, B. L. 2019, ApJ, 876, 155
-- Sampson et al. (2024) Sampson, M. L., Melchior, P., Ward, C., & Birmingham, S. 2024, Astronomy and Computing, 49, 100875
-- Sargsyan et al. (2023) Sargsyan, A., Navasardyan, S., Xu, X., & Shi, H. 2023, in IEEE/CVF International Conference on Computer Vision (IEEE), 7335–7345
-- Smith et al. (2022) Smith, M. J., Geach, J. E., Jackson, R. A., et al. 2022, MNRAS, 511, 1808
-- Spagnoletti et al. (2024) Spagnoletti, A., Boucaud, A., Huertas-Company, M., Kabalan, W., & Biswas, B. 2024, arXiv preprint arXiv:2411.19158
-- Stone & Courteau (2019) Stone, C. & Courteau, S. 2019, ApJ, 882, 6
-- Stone et al. (2021) Stone, C., Courteau, S., & Arora, N. 2021, ApJ, 912, 41
-- Storey-Fisher et al. (2024) Storey-Fisher, K., Hogg, D. W., Rix, H.-W., et al. 2024, ApJ, 964, 69
-- Thanh-Tung & Tran (2020) Thanh-Tung, H. & Tran, T. 2020, in 2020 international joint conference on neural networks (ijcnn), 1–10
-- Wang et al. (2018) Wang, Y., Tao, X., Qi, X., Shen, X., & Jia, J. 2018, Advances in neural information processing systems, 31
-- Yeh et al. (2017) Yeh, R. A., Chen, C., Yian Lim, T., et al. 2017, in IEEE conference on computer vision and pattern recognition (IEEE), 5485–5493
-- Yu et al. (2018) Yu, J., Lin, Z., Yang, J., et al. 2018, in IEEE conference on computer vision and pattern recognition (IEEE), 5505–5514
-- Zhang et al. (2023) Zhang, L., Rao, A., & Agrawala, M. 2023, in IEEE/CVF International Conference on Computer Vision (IEEE), 3836–3847
-Appendix A Reconstruction metrics
-Figures Fig. 19 show the differences in the rate of change measurement between input images and the inpainting output.
-Appendix B Extreme examples
+Appendix A Masks
+To explore the ability of the model at recreating galaxy images, we apply a variety of different masks to a selection of unseen images.
+Figure 17 shows the original image on the leftmost column, followed by the regenerated images in subsequent columns.
+With each mask shown on the top row, the white area shows the pixels that are fixed from the original image, whereas the black areas are generated using the diffusion model.
+As our model’s aim is not to generate large collections of synthetic data, the variety and scientific validity of the produced galaxies are not put under high scrutiny.
+Instead, we only check that there are no obvious catastrophic mistakes.
+In the last two columns, we can see that due to the large mask and fewer pixels to guide the reconstruction, some images produce a visible boundary around the image where the combination of fixed and generated pixels is not as seamless as the other images.
+Given that the masks used for AGN detection are significantly smaller than those presented in Fig.
+17, the fraction of pixels that the inpainting is conditioned on will be much higher, providing tighter constraints on the output compared to these examples.
+Improvements to the scheduler, as discussed in Sect.
+6.1, would also reduce such artifacts.
+Appendix B Alternative loss
+Making use of standard metrics whose primary use-case is data on a consistent scale, such as 0–255 pixels, may not function optimally with the high dynamic range of VIS data.
+Specifically, since MSE is not scale invariant, having images that feature pixels that are an order of magnitude brighter, results in relative errors increasing by a square of that difference.
+This is typically a benefit as it forces outlier predictions to be severely penalised.
+However, with our data, this results in brighter images being disproportionately penalised compared to fainter ones, even if the predictions are both the same relative distance away from the true prediction.
+As shown in Sect.
+2.3, we applied a normalisation to by dividing each image’s error by its respective maximum pixel value, which could be from a brighter external source.
+To visualise the impact of amending the loss function, with respect to an images brightness, Fig.
+18 shows how the error from the unmodified MSE changes as the model is trained for longer.
+Each line represents the error for a collection of images in a random batch, with the increased transparency indicating the passage of time where the model had been trained on more epochs.
+The final epoch of each model is presented by a black line.
+The effect of removing point-like sources, as discussed in Sect.
+4.1, is also explored.
+Although all four models arrive at a similar minimum loss, and begin the training process near-identically, their respective behaviour throughout training differs significantly.
+For the full data set with the default loss (top left of Fig.
+18), training prioritises brighter images, leading to some significantly low errors towards the rightmost data.
+However, this comes at the expense of fainter data, where wild fluctuations in error show that the model’s output is not very robust when it comes to precise predictions.
+This happens even when MSE might suggest low error, since pixel values below 1.0 result in minimal losses despite the predictions being relatively incorrect.
+For the normalised loss with the full data set (top right), since the penalisation is shared more fairly across pixel values, it is the density of images that has the most influence.
+Early on in training we can see that as the majority of data reduces in error, the brightest images remain at near-starting performance.
+This continues until the errors of the fainter images become so low that model is forced to prioritise the bright images.
+The key difference here is this only happens after the rest of the images are at a sufficient standard.
+The final outcome is fainter images performing well and the entire scale of images having the same maximum error.
+The two models which feature the reduced training set (bottom row) share similar dynamics during training, regardless of the loss used.
+This is largely due to many point-like sources also being some of the brightest images and therefore the scale of the data set is reduced.
+As the most problematic data are removed, the impact of disproportionate penalisation is significantly reduced, allowing the fainter sources to be the focus during training.
+Appendix C Reconstruction metrics
+Figure Fig. 19 show the differences in the rate of change measurement between input images and the inpainting output.
+Appendix D Extreme examples
 We present output examples from the inpainting pipeline where pixel values deviate significantly from the expected input pixels.
 Figure 20 highlights much fainter pixels, while Figure 21 shows much brighter ones.
-Appendix C Morphology examples
+Appendix E Morphology examples
 We show output examples from various morphology types to showcase how the diffusion model adapts to different galaxy shapes.
@@
 24, respectively.
-Appendix D Additional results
+Appendix F Additional results
 We present collated results for each of the star, galaxy, AGN and QSO selections in Fig.